### v2

# Paper2Code: Automating Code Generation from Scientific Papers in Machine Learning

Abstract
Despite the rapid growth of machine learning research, corresponding code implementations are often unavailable, making it slow and labor-intensive for researchers to reproduce results and build upon prior work. In the meantime, recent Large Language Models (LLMs) excel at understanding scientific documents and generating high-quality code. Inspired by this, we introduce PaperCoder, a multi-agent LLM framework that transforms machine learning papers into functional code repositories. PaperCoder operates in three stages: planning, where it constructs a high-level roadmap, designs the system architecture with diagrams, identifies file dependencies, and generates configuration files; analysis, which focuses on interpreting implementation-specific details; and generation, where modular, dependency-aware code is produced. Moreover, each phase is instantiated through a set of specialized agents designed to collaborate effectively across the pipeline. We then evaluate PaperCoder on generating code implementations from machine learning papers based on both model-based and human evaluations, specifically from the original paper authors, with author-released repositories as ground truth if available. Our results demonstrate the effectiveness of PaperCoder in creating high-quality, faithful implementations. Furthermore, it consistently shows strengths in the recently released PaperBench benchmark, surpassing strong baselines by substantial margins. Code is available at: https://github.com/going-doer/Paper2Code.
1 Introduction
Reproducibility lies at the heart of scientific progress, which enables researchers to validate findings, build upon prior work, and ultimately push the boundaries of knowledge [7, 3, 32]. However, across many disciplines, reproducing scientific results remains an enduring challenge. This is often due to incomplete documentation, missing experimental details, lack of access to data or proprietary tools, and, especially in machine learning research, the absence of corresponding code: for example, only 21.23 % of the papers accepted to top-tier machine learning conferences in 2024 provide their code implementations shown in Figure 1. As a result, researchers frequently invest substantial effort in reverse-engineering methods and experimental results from papers, a process that is both time-consuming and labor-intensive, subsequently slowing down the overall pace of scientific innovation.
Meanwhile, recent Large Language Models (LLMs) have shown outstanding capabilities in understanding and generating both natural language and programming code [12, 30, 37], with performances increasingly approaching or even surpassing that of domain experts in some scenarios. In addition, this progress has sparked growing interest in leveraging LLMs to accelerate scientific workflows, particularly in the early stages of ideation for new and valid research hypotheses [26, 46, 38, 2, 20, 47, 39]. Furthermore, some of these studies, as well as others focusing on later stages of automating experimental validations and improvements [16, 49, 4, 42], demonstrate the potential of LLMs to generate code and even carry out experiments end-to-end; however, they typically assume and heavily rely on access to pre-existing implementations, partial code snippets, or well-defined APIs. As such, it remains questionable whether generating complete and faithful implementations solely from research papers (without access to prior code, APIs, or additional supplementary materials) can be achievable.
To answer this question, we introduce PaperCoder, a multi-agent LLM-powered framework, designed to automatically generate executable code repositories in machine learning directly from and contextualized with research papers, which differs from prior work that requires partial implementations from human inputs. Specifically, PaperCoder aims to emulate the typical life cycle of human developers and researchers in writing the repository-level code, by decomposing the task into three structured stages: planning, analysis, and generation. First, during the planning stage, the proposed framework constructs a high-level roadmap to identify core components to implement, draws the overall system architecture with class and sequence diagrams to model structural relationships between modules, identifies file dependencies with their execution orders to guide correct build and execution flows, and generates configuration files to enable flexible customization of experimental workflows by human researchers. This is followed by the analysis stage, performing a fine-grained interpretation of each file and function with respect to their intended functionality, such as required inputs and outputs, interactions with other modules, and any algorithmic or architectural constraints derived from the source paper. Finally, in the generation stage, the framework synthesizes the entire code base based on the execution order determined earlier, along with the artifacts produced in the previous stages.
(a) PaperCoder overview
(b) Code availability
Figure 1: (a) PaperCoder overview. The proposed PaperCoder aims to transform given scientific papers (in machine learning domains) into code repositories, which consists of three sequential steps: planning, analyzing, and coding. (b) Code availability. The availability of the code repositories, where blue bars indicate the total number of accepted papers and orange regions represent the subset of papers with officially released code.
To validate the effectiveness of PaperCoder, we conduct extensive evaluations on a subset of recent machine learning papers accepted at top-tier venues in 2024, including NeurIPS, ICML, and ICLR referred to as our proposed Paper2Code benchmark. Furthermore, we also incorporate the recently released PaperBench benchmark in our evaluation suite, enabling the fine-grained evaluations of code implementations with some papers from ICML 2024 [40]. Then, on a battery of tests conducted with automated model-based evaluations (covering both reference-free and reference-based settings, conditional on the availability of author-released ground-truth repositories) as well as expert human evaluations (based on authors of original papers), PaperCoder demonstrates substantial improvements over baselines, generating more valid and faithful code bases that could meaningfully support human researchers in understanding and reproducing prior work. Specifically, 77% of the generated repositories by PaperCoder are rated as the best, and 85% of human judges report that the generated repositories are indeed helpful. Also, further analyses show that each component of PaperCoder (consisting of planning, analysis, and generation) contributes to the performance gains, but also that the generated code bases can be executed, sometimes with only minor modifications (averaging 0.48% of total code lines) in cases where execution errors occur.
2 Related Work
Large Language Models with Code
Large language models (LLMs) demonstrate impressive capabilities in both text understanding and generation [30, 12, 37]. Recently, their applicability has extended beyond general knowledge tasks into specialized fields such as mathematics, science, and coding, showcasing impressive reasoning and knowledge representation capabilities [34, 43, 41]. Particularly, code-specialized LLMs [10, 9, 17] have garnered significant attention by achieving remarkable performance on various software engineering tasks [45], including software development, software design [36, 15], requirements elicitation [29], and specification formalization [27]. Our work aligns closely with this line of research, exploring and expanding upon the capabilities and applications of code-specialized LLMs.
Repository-Level Coding
LLM-based code generation can be broadly categorized into single-file and multi-file (repository-level) coding. Single-file coding focuses on generating relatively short code snippets to solve isolated tasks, such as programming competition problems or simple coding queries, and was the primary focus in the early stages of LLM-based coding research [21, 6, 1, 14]. As LLMs have advanced in code comprehension, long-context reasoning, and handling complex workflows, research has increasingly shifted toward repository-level coding, where multiple files are generated while jointly considering architectural and functional requirements [23, 18, 25]. Several recent studies have explored this direction [48, 31]. Multi-agent frameworks such as ChatDev leverage LLMs to collaborate through structured dialogues [36], while MetaGPT adopts a waterfall development model with role-specific generation stages [15]. These advancements highlight the growing capability of LLMs to support end-to-end software engineering tasks beyond single-file generation.
LLM-Powered Research
Scientific progress typically involves a cycle of idea generation, hypothesis validation, and experimental execution [33]. LLMs have been applied to various stages of this cycle, including research ideation [2, 20], hypothesis generation [47, 35], and peer review [8, 22, 44], thereby helping researchers overcome existing limitations and accelerate scientific discovery [19, 26, 46]. In computer science and machine learning, where code-based experimentation is fundamental, LLMs have also been utilized to design experiments that enhance existing codebases. Many recent studies focus on refining implementations or optimizing performance, assuming the availability of the original implementation [16, 49, 4, 42]. However, this assumption limits their applicability in cases where source code is unavailable. Reproducibility plays a critical role in scientific discovery, enabling researchers to identify limitations, refine methodologies, and build upon prior work [32]. To evaluate LLMs’ replication capabilities, PaperBench introduced a benchmark in which AI agents attempt to reproduce machine learning papers [40]. While this benchmark emphasizes evaluation, our work addresses a challenge of automatically generating code repositories from scientific papers in the absence of original implementations. To this end, we propose a framework that leverages LLMs for end-to-end research reproducibility. By focusing on repository-level reproduction, our approach expands the scope of LLM-powered automation beyond ideation and hypothesis generation, contributing to a critical yet underexplored aspect of scientific research.
3 Methods
3.1 Problem Definition
Research Repository Generation
In machine learning research, experiments are typically conducted using a code. However, in many cases, researchers do not release their code, making it challenging for others to reproduce and verify the proposed method and experiments [28, 32]. When non-authors attempt to re-implement the code manually, the process is often labor-intensive, time-consuming. To address these challenges, we define the task of Research Repository Generation. The goal of this task is to read a paper and generate a corresponding repository that implements its methods and experiments. We define a repository as a collection of one or more code files necessary for reproducing the methods and experiments described in the paper. By automating this process, research repository generation facilitates the validation and reproducibility of scientific findings. Additionally, new research ideas often build upon prior work. However, when an researcher does not release their code, experimental validation becomes difficult, potentially hindering further advancements. Our approach aims to bridge this gap by enabling researchers to experimentally verify prior work, even when the original implementation is unavailable, thus contributing to a more reproducible and transparent research. In this study, we frame repository generation as a software development problem and assume that a model is responsible for writing the code to create a functional software repository. Given a research paper as input, the model generates a repository containing the code necessary to replicate the paper’s methods and experiments. Formally, we define this as: $M(R)=C$; where $R$ represents the paper, $C$ represents the generated code, and $M$ is a function or model. While $M$ can take various forms, large language models (LLMs) are a representative candidate. A language model functions as a mapping from an input $x$ (text) to an output $y$ (text or code), making it well-suited for this task.
Figure 2: (Left) The naive approach, where a model directly generates code from the paper. (Right) PaperCoder framework, which decomposes the task into three stages: (1) Planning, where a high-level implementation plan is constructed based on the paper’s content, including overall plan, architectural design, logic design, and configuration files; (2) Analyzing, where the plan is translated into detailed file-level specifications; and (3) Coding, where the final codes are generated to implement the paper’s methods and experiments.
3.2 Our Method
We introduce PaperCoder, a novel framework designed specifically for research repository generation that implements the methods and experiments described in a research paper, illustrated in Figure 2. The goal of our work is to model the repository generation process as: $M(R)=C$ where $M$ is the model responsible for transforming the paper into a code repository, $R$ represents the research paper, and $C$ represents the generated code. Inspired by software development methodologies, we adopt a structured approach that mirrors well-established, empirically validated software engineering principles. Specifically, we follow a plan-analysis-implementation workflow to systematically address the problem of research repository generation. To achieve this, we decompose the process into three sequential phases: 1) Planning. 2) Analyzing. 3) Coding. Each phase utilizes task-specialized LLM agents that leveraging a multi-agent scenarios. Formally, we define the repository generation process as: $C=M(P)=M_{\text{code}}(R,P,A)$; where $R$ represents the research paper, $P$ is the plan and $A$ indicates the analysis. Each components are generated as follows: $P=M_{\text{plan}}(R)$, $A=M_{\text{analysis}}(R,P)$ and $C=M_{\text{code}}(R,P,A)$. The following sections provide a detailed explanation of each phase.
3.2.1 Planning
Providing the research paper as a input to the model and expecting it to generate a complete repository is highly challenging. Research papers are primarily written to document discoveries and persuade readers, rather than to serve as structured input for software development. Consequently, papers often contain supplementary information that, while essential for conveying core concepts, is not directly relevant to implementation. It could be introduces noise, making repository generation difficult and less effective. To address this challenge, instead of solely using the paper as input, we propose decomposing the paper into a structured multi-aspect plan. This approach organizes the key implementation-relevant elements into four distinct components, ensuring that the generated repository is well-structured and aligned with the paper’s methodology. Formally, $M_{\text{plan}}(R)=P=\{o,d,l,g\}$; where $o$ indicates the overall plan, $d$ is the architecture design, $l$ represents the logic design and $g$ is the configuration file. Each output is generated in each phase, and this phase follows a sequential structure, where the output of each stage serves as the input for the next. Note that each phase utilized by the same agent which is the plan agent $M_{\text{plan}}$.
Overall Plan
The first step within planning phase, the overall plan, involves summarizing and organizing the core elements necessary to implement the research repository from a high-level perspective. This summary provides an essential conceptual framework that guides subsequent steps clearly. Formally, $M_{\text{plan}}(R)=o$, where $R$ is the research paper and $o$ indicates the overall plan.
Architecture Design
The second stage involves architecting the software based on the overall plan generated in the previous stage and the research paper. Designing a well-structured architecture is essential, particularly for software systems where multiple features must interact seamlessly. This stage focuses on identifying the necessary components and defining their relationships to ensure a well-organized and functional repository. To achieve this, we asks to create key artifacts that define the software architecture. The file list is a structured collection of files required for the repository, outlining the software’s modular structure. The class diagram provides a static representation of the system’s data structures and interfaces. Using Unified Modeling Language (UML) notation, a standardized visual language for modeling software systems, we depict classes as rectangles, attributes and methods as lists, and relationships as connecting lines to illustrate how different components interact. The sequence diagram represents the program’s call flow and object interactions dynamically, which uses UML notation to show actors as objects, messages as arrows, and lifelines as dashed lines, visually mapping out how components communicate over time. This structured approach ensures a clear and organized representation of the software’s architecture. By constructing these artifacts, we visually represent the essential components described in the research paper, enabling a more structured and systematic approach to repository generation. This process facilitates better analysis of dependencies and relationships, ensuring that the generated repository aligns with the paper’s core idea. Formally, we define the architecture design process as: $M_{\text{plan}}(R,o)=d$ where $R$ represents the paper, $o$ refers to the previously generated overall plan, and $d$ denotes the resulting architectural design artifacts.
Logic Design
In software development, individual files rarely function in isolation. Instead, they generally exhibit interdependencies through imports and module interactions. For example, if a function A is defined in utils.py and later imported into evaluation.py, utils.py must be implemented before evaluation.py to maintain a correct dependency structure. To account for these dependencies, this stage takes as input the research paper along with the artifacts generated in the previous two stages. It then analyzes the logic of each file and its components, determining the necessary dependencies and the optimal execution order for implementation. As an output, it produces an ordered file list, detailing each file’s role which files should be implemented considering dependencies and its dependencies within the repository. This approach ensures that repository generation considers not only individual file structures but also inter-file communication, facilitating a well-organized and logically coherent implementation. Formally, we define this dependency-aware file ordering process as $M_{\text{plan}}(R,o,d)=l$ where $R$ represents the research paper, $o$ refers to the overall plan, $d$ denotes the architectural specifications, and $l$ is the ordered file list which is the structured output defining the order and role of each file in the repository.
Configuration File Generation
Finally, the configuration file generation step synthesizes all previously established outputs to produce a configuration file (config.yaml) containing hyperparameters and configuration required for model training. This process takes as input the overall plan, architectural design, and ordered file list produced in the previous stages. At this stage, users can review and modify the config.yaml file to identify and correct any missing or incorrectly specified details. For example, users may need to specify the path to a Hugging Face dataset or define the checkpoint storage directory. This step helps reduce hallucinations during the generation process, such as the model producing nonexistent datasets or referencing incorrect file paths. Formally, the configuration generation process is defined as $M_{\text{plan}}(R,o,d,l)=g$ where $R$ represents the research paper, $o$ refers to the overall plan, $d$ denotes the software architecture specifications, $l$ represents the logic design, and $g$ is the resulting configuration file.
3.2.2 Analyzing
While the planning phase primarily focuses on designing the overall repository structure and outlining a high-level roadmap, the analyzing phase delves into the implementation specifics of each individual file. In this phase, the detailed purpose and necessary considerations for each file within the repository are thoroughly analyzed. The outputs generated at this stage clearly specify what each file should achieve and highlight critical factors required for successful implementation. Specifically, the inputs to this analysis phase consist of the original research paper and previously generated artifacts (the overall plan, architecture design, logic designs, and configuration file). The outputs of this phase include file-level analyses documenting precise implementation details, which will later inform the code generation process. Formally, the file-level analysis process is defined as $\{M_{\text{analysis}}(R,P=\{o,d,l,g\},f_{i})=a_{i}\}_{i=1}^{n=|\text{file}|};\text{where}f\in F,a\in A$ where $R$ represents the research paper, $P$ includes four plans, which $o$ refers to the overall plan, $d$ denotes the architectural specifications, $l$ represents the logic design which consists of structured dependencies among files, $g$ is the configuration file, and $f_{i}$ specifies the target file for which the analysis is being conducted. This process is applied to each file individually to generate a $a_{i}$ which is the comprehensive file-level implementation plan.
3.2.3 Coding
The final stage of our methodology is the coding phase, which produces code constituting the research repository. The generation of each file is guided by the comprehensive outputs from previous phases: the research paper itself, the overall plan, architecture design, logic designs, the configuration file, file-specific analyses, and previously generated code. As repository files often exhibit import dependencies among themselves, we adhere strictly to the ordered file list established during the planning phase, ensuring sequential consistency. Formally, the file-specific code generation process is defined as $\{M_{\text{coder}}(R,P=\{o,d,l,g\},f_{i},a_{i})=c_{i}\}_{i=1}^{n=|\text{file}|};\text{where }f\in F,a\in A,c\in C$. This process is applied sequentially for each file to ensure a structured and logically coherent repository implementation $c_{i}$. The initially produced code may require subsequent debugging or refinement to ensure correctness and full functionality. In this work, comprehensive debugging strategies and detailed error-correction workflows remain beyond the current scope of this paper.
4 Experimental Setup
4.1 Data
Paper2Code Benchmark
To evaluate the effectiveness of our framework for reproducing both the proposed methods and experiments described in scientific papers, we constructed a experimental benchmark based on papers from ICML 2024, NeurIPS 2024, and ICLR 2024. Using the OpenReview API [^1] https://docs.openreview.net/reference/api-v2, we first filtered the accepted papers from each conference to include only those with publicly available GitHub repositories. Among these, we selected repositories with a total codebase of fewer than 70,000 tokens to ensure reproducibility within a manageable scope. To ensure quality, we performed model-based evaluations of the repositories and selected the top 30 highest-scoring papers from each venue. These 90 papers constitute our Paper2Code benchmark for experimental validation. A detailed list of the 90 papers is provided in Table 9, Table 10 and Table 11. For the human evaluation, we select 13 papers authored by the evaluation participants, and listed in Table 13.
PaperBench Code-Dev
We further validate our framework using the PaperBench Code-Dev benchmark [40], which consists of a curated set of 20 papers from ICML 2024.
4.2 Baselines and Our Model
Paper2Code Benchmark
To the best of our knowledge, there is no existing framework that directly addresses the task of end-to-end paper-to-code reproduction. However, we consider several related approaches as baselines, particularly those designed for software development from natural language inputs such as requirements. These serve as the closest comparison points in terms of producing functional code repositories. In addition, we include comparisons with a number of ablated variants that serve as simple or naive baselines, as follow: 1. ChatDev [36] - It is a multi-agent framework where LLM-driven agents collaboratively develop software through dialogues. In our experiments, the entire paper is provided as input(requirements) to generate a full code repository; 2. MetaGPT [15] - It adopts a role-based multi-agent paradigm, where software development is organized via Standardized Operating Procedures (SOPs), which specify structured workflows and responsibilities for each agent. We provide the entire paper as input and prompt the system to construct the full code repository; 3. Abstract - A naive baseline where only the paper’s abstract is provided to the language model, requiring it to implement the code repository based on minimal information; 4. Paper - Another naive baseline in which the full paper is provided as input, and the model is prompted to generate a corresponding code repository; 5. PaperCoder (Ours) - Our proposed method, which integrates planning, analysis and code generation to generate high-quality repositories from scientific papers; 6. Oracle - The official code repository released by the paper’s authors. This serves as an upper bound for comparison, reflecting the intended implementation of the proposed in the paper.
PaperBench Code-Dev
We compare our method against the two agent variants reported in the original benchmark. The Basic Agent follows a ReAct-style approach and is equipped with a predefined set of tools, including a bash shell command executor, a Python code runner, a web browser, and a paginated file reader for processing long documents. The Iterative Agent extends the Basic Agent by utilizing the full available computation time and incorporating custom prompting strategies that encourage step-by-step reasoning through subgoal decomposition (see Starace et al. [40] for details).
4.3 Evaluation Setups
4.3.1 Paper2Code Benchmark
Evaluating code generation typically relies on unit tests or verifying specific implementation components [5, 11]. However, such evaluations are often infeasible due to the lack of gold-standard repositories. Many papers do not release official code, and even when they do, test scripts are frequently unavailable. Additionally, manually annotating each paper with ground-truth implementation details is highly labor-intensive, rendering large-scale evaluation impractical.
Model-Based Evaluation
To address this challenge, we adopt a model-based evaluation approach [50, 13, 24] that assesses repository quality with or without access to gold-standard repository. We consider two variants: (1) reference-based, which leverages both the paper and the author-provided repository, and (2) reference-free, which evaluates the generated repository using only the paper. For both variants, we prompt the language model to critique the required implementation components and evaluate their correctness. It assigns severity levels (high, medium and low) to any missing or flawed elements and generates an overall correctness score on a 1 to 5 scale. To ensure stability, we report the average score across multiple generations using $n$-way sampling. In this experiments, we set the $n=8$ and use the o3-mini-high as a evaluation model.
Reference-Based Evaluation
When an official repository is available, we treat it as one plausible implementation of the method described in the paper. While multiple valid implementations may exist, the author-released version is generally considered the most accurate, as it best captures the core components necessary for reproducing experimental results. In this setting, we provide both the paper and the gold-standard repository as input, and prompt the evaluation model to identify and critique the required components. The model then compares the predicted repository against these components and assigns a correctness score from 1 to 5, reflecting both component coverage and the severity of any errors.
Reference-Free Evaluation
In many cases, author-released code is unavailable. To handle such scenarios, we propose a reference-free evaluation strategy that relies solely on the paper and the generated repository. As in the reference-based setting, the model is prompted to infer the required implementation components from the paper, critique them, and assess whether they are adequately implemented in the generated repository. A correctness score from 1 to 5 is then assigned based on the presence and quality of these components.
Human Evaluation
[…]
4.3.2 PaperBench Code-Dev Evaluation
For evaluating our system on the PaperBench Code-Dev benchmark, we adopt the official evaluation protocol, which measures replication accuracy across a curated set of ICML 2024 papers. Specifically, we follow the rubric authored by the original paper’s authors, which defines a hierarchical set of implementation requirements. The evaluation model assesses each generated repository based on whether the submitted code correctly implements any of the specified requirements. Importantly, the evaluation focuses only on the code development node which whether the candidate repository contains a correct implementation of some requirement.
Table 1: Results on our experimental Paper2Code benchmark. We report average model-based evaluation scores for each conference. Oracle denotes the official repository released by the paper’s authors. Reference-based evaluation assesses correctness (on a 1–5 scale) by comparing the generated repository against both the paper and the official implementation, while reference-free evaluation relies solely on the paper. We report statistics on the average number of tokens, files, and functions per repository. The best scores are highlighted in bold.
Reference-based Reference-free Statistics
ICML NeurIPS ICLR ICML NeurIPS ICLR # of Tokens # of Files # of Funcs
ChatDEV 2.97 (0.58) 2.96 (0.69) 2.70(0.63) 4.12 (0.53) 4.01 (0.74) 4.00 (0.65) 6150.54 6.99 23.82
MetaGPT 2.75 (0.70) 2.95 (0.87) 2.48 (0.48) 3.63 (0.75) 3.59 (0.92) 3.52 (0.60) 5405.21 3.24 18.08
Abstract 2.43 (0.49) 2.35 (0.62) 2.28 (0.42) 3.01 (0.60) 2.99 (0.78) 3.03 (0.64) 3376.99 1.28 12.62
Paper 3.28 (0.67) 3.22 (0.80) 3.08 (0.66) 4.30 (0.53) 4.08 (0.84) 4.15 (0.63) 3846.33 1.79 14.84
PaperCoder 3.72 (0.54) 3.83 (0.50) 3.68 (0.52) 4.73 (0.44) 4.77 (0.38) 4.73 (0.32) 14343.38 6.97 35.22
Oracle - - - 4.80 (0.32) 4.83 (0.38) 4.84 (0.26) 32149.04 28.00 122.03
5 Experimental Results and Analysis
Main Results
Table 1 presents the main experimental results, our method consistently outperform across all conferences and both evaluation modes. Under the reference-based setting, our approach achieves the highest average correctness scores of 3.72, 3.83, and 3.68 on ICML, NeurIPS, and ICLR papers, respectively. Similarly, in the reference-free evaluation, our method attains scores of 4.73, 4.77, and 4.73, outperforming all baselines. Compared to software development work baselines such as ChatDev and MetaGPT, our method exhibits substantial performance gains. Notably, although ChatDev generates a comparable number of files (6.99 vs. 6.97), our method produces significantly more functions (35.22 vs. 23.82), indicating a higher level of granularity and completeness in the generated repositories. In contrast, MetaGPT lags behind both in evaluation scores and in code quantity metrics. Ablation variants using only the abstract or full paper underperform relative to ours. This highlights the effectiveness of our multi-stage framework in systematically extracting and organizing implementation-relevant information. For reference, we include an Oracle score that represents the upper bound, corresponding to the official implementation by the authors. While there is still a gap to the Oracle, our method achieves comparable results that are closer to this upper bound compared to existing LLM-based frameworks. Overall, these results demonstrate that our proposed system not only surpasses prior approaches in generating accurate and comprehensive code repositories but also excels in producing fine-grained, detail-oriented implementations necessary for faithful reproduction of research methods.
Figure 3: Correlation between reference-based and reference-free model-based evaluations.
Correlation between Reference-based and Reference-free
We investigate the alignment between reference-based and reference-free evaluation metrics by computing their correlation across a large set of generated repositories. As shown in Figure 3, the two evaluation scores exhibit a strong positive correlation, with a Pearson correlation coefficient of $r=0.79$ and a statistically significant $p$-value of $0.00$. This high correlation suggests that reference-free evaluation which does not rely on access to ground-truth code can serve as a reliable proxy for reference-based assessment. The result supports the feasibility of using reference-free methods in scenarios where official implementations are unavailable, while still maintaining consistency with traditional evaluation standards.
Table 2: Results on the PaperBench Code-Dev benchmark using o3-mini-high. For BasicAgent and IterativeAgent, results are averaged over three runs, and standard errors are reported.
Model Replication scores (%)
BasicAgent $5.1\pm 0.8$
IterativeAgent $16.4\pm 1.4$
PaperCoder 44.26
PaperBench Code-Dev Results
To assess the effectiveness of our approach in realistic research scenarios, we evaluate it on the PaperBench Code-Dev benchmark. As shown in Table 2, our method achieves a best replication score. This demonstrates the superiority of our structured multi-agent pipeline, which performs coordinated planning, analysis, and implementation.
Table 3: Human evaluation results. For model-based evaluations, including reference-based and reference-free settings, correctness scores are converted into rankings for comparability. In the human evaluation, 13 authors were asked to rank repositories within two groups based on method type. Rankings are then converted into scores by assigning 5, 3, and 1 points to the first-, second-, and third-ranked repositories, respectively.
Score ($\uparrow$) Ranking ($\downarrow$)
Ref-based Ref-free Human Ref-based Ref-free Human
Abstract 2.36 (0.34) 2.99 (0.53) 1.62 (1.26) 3.00 (0.00) 3.00 (0.00) 2.69 (0.63)
Paper 3.19 (0.44) 4.22(0.52) 3.15 (1.28) 1.86 (0.36) 1.79 (0.43) 1.92 (0.64)
PaperCoder (Ours) 3.74 (0.30) 4.71 (0.27) 4.23 (1.30) 1.14 (0.36) 1.07 (0.27) 1.38 (0.65)
ChatDev 2.67(0.63) 3.87 (0.36) 2.69 (1.11) 2.50 (0.52) 2.36 (0.50) 2.15 (0.55)
MetaGPT 2.65 (0.46) 3.38 (0.66) 1.77 (1.30) 2.07 (0.52) 2.14 (0.52) 2.61 (0.65)
PaperCoder (Ours) 3.74 (0.30) 4.71 (0.27) 4.54 (1.20) 1.00 (0.00) 1.00 (0.00) 1.23 (0.60)
Human Evaluation Results
To assess the practical utility and subjective quality of the generated code, we conduct a comprehensive human evaluation. We include both model-based evaluations (reference-based and reference-free) and direct human assessments, where authors or expert annotators rank the quality of each generated repository. Table 3 summarizes the results across these three evaluation types. Across all evaluation criteria, our method consistently achieves the highest scores and the best rankings. In terms of average human evaluation scores, our system achieves 4.23 and 4.38 comparing both ablation variations and software development works, respectively, substantially outperforming all baselines. Similarly, in model-based reference-based and reference-free evaluations, our method attains 3.74 (ref-based) and 4.71 (ref-free), outperforming alternatives such as ChatDEV and MetaGPT. In the ranking-based evaluation, our method also secures the top position across all metrics. Notably, it achieves the best average rank in human evaluations (1.38 and 1.31), while other baselines such as MetaGPT and only Abstract consistently rank lower. The low variance in our ranking (e.g., 1.00 in both reference-based and reference-free) further demonstrates the robustness of our system across diverse evaluation settings. These results clearly indicate that our proposed approach produces more useful and accurate implementations than existing methods, regardless of whether the evaluation is automated or conducted by human experts.
Table 4: Human evaluation results on model variants of PaperCoder. DS-Coder refers to DeepSeek-Coder-V2-Lite-Instruct, Qwen-Coder to Qwen2.5-Coder-7B-Instruct, and DS-Distil-Qwen to DeepSeek-R1-Distill-Qwen-14B. o3-mini-high denotes the high reasoning-effort of o3-mini. Other settings follow Table  3.
DS-Coder Qwen-Coder DS-Distill-Qwen o3-mini-high
Score ($\uparrow$) Ref-based 1.63 (0.43) 1.8 (0.28) 2.07 (0.30) 3.74 (0.30)
Ref-free 1.82 (0.39) 2.1 (0.28) 2.29 (0.29) 4.71 (0.27)
Human 1.41 (0.64) 2.74 (1.14) 3.05 (1.04) 4.79 (0.74)
Ranking ($\downarrow$) Ref-based 3.36 (0.93) 2.93 (0.62) 2.36 (0.63) 1.00 (0.00)
Ref-free 3.36 (0.84) 2.86 (0.66) 2.14 (0.36) 1.00 (0.00)
Human 3.69 (0.48) 2.69 (0.85) 2.46 (0.78) 1.15 (0.55)
Detailed Analysis on Generated Repository
We conduct a fine-grained human evaluation to better understand the practical utility and implementation quality of the generated repositories. Specifically, we asked authors to review a set of repositories about covering both baselines and ablated variants and to select the one they considered most suitable for reproducing their work. 77% (10 out of 13 authors) selected the repository generated by our method as the top choice, followed by 3 authors who preferred the Paper variant. The most common reasons cited for favoring our method include completeness, clean structure, and faithfulness to the original paper. Full results are shown in Table 8.
Figure 4: Human evaluation results on the helpfulness and section-level implementation accuracy of repositories generated by our framework.
[…]
Table 5: Rank correlation between model-based and human evaluations is measured using GPT-4o (gpt-4o-2024-11-20) and o3-mini with high reasoning-effort.
Reference-based Reference-free
GPT-4o o3-mini-high GPT-4o o3-mini-high
0.67 0.71 0.67 0.67
Analysis on Human Alignment for Evaluation
We measure their correlation with human judgments for assessing the reliability of model-based evaluation metrics. Table 5 reports the rank correlation between human scores and model-based scores across both reference-based and reference-free settings, using two evaluation models: GPT-4o and o3-mini-high. The results show moderate to strong correlations in all settings. In the reference-based evaluation, o3-mini-high achieves the highest correlation with human judgments (0.67), followed closely by GPT-4o (0.63). In the reference-free setting, both models exhibit strong alignment with human scores, each achieving a correlation of 0.63. These results suggest that model-based evaluations can serve as reliable proxies for human judgment. This supports the use of automatic scoring in large-scale experiments and reinforces the credibility of the evaluation results presented in this study. Based on these findings, we adopt the o3-mini-high model as the evaluation model in all experiments.
Analysis on Different LLMs
To investigate the impact of the underlying language model on overall performance, we experiment with four different LLM backbones: DS-Coder (DeepSeek-Coder-V2-Lite-Instruct), Qwen-Coder (Qwen2.5-Coder-7B-Instruct), DS-Distill-Qwen (DeepSeek-R1-Distill-Qwen-14B), and o3-mini-high (the high reasoning-effort variant of o3-mini). Each model is used within the our framework, and we evaluate the outputs using reference-based, reference-free, and human evaluation metrics. Results are summarized in Table 4. We observe that o3-mini-high consistently outperforms all other models across all evaluation dimensions. In terms of reference-based and reference-free scores, it achieves 3.74 and 4.71, respectively, with a higher score of 4.79 in human evaluation. Notably, it positions the top ranking in all settings, with an average rank of 1.00 in both reference-based and reference-free metrics, and 1.15 in human evaluations. Among the other models, DS-Distill-Qwen performs best, followed by Qwen-Coder and DS-Coder. This performance trend holds consistently across both model-based scores and human judgments. These results highlight the importance of selecting a capable and well-aligned backbone model for generating high-quality, executable repositories from research papers. Based on this results, we mainly use the o3-mini-high as the basis backbone model for all experiments.
Table 6: Ablation study results on PaperCoder. We report the average scores and standard deviations on the ICML 2024 subset of Paper2Code benchmark using model-based reference-based and reference-free evaluation.
Ref-based Ref-free
Abstract 2.43 (0.49) 3.01 (0.60)
Paper 3.28 (0.67) 4.30 (0.53)
+ Overall Plan 3.40 (0.57) 4.34 (0.58)
+ Arch. Design 3.13 (0.68) 4.07 (0.74)
+ Logic Design 3.60 (0.52) 4.50 (0.57)
+ Config File 3.66 (0.45) 4.45 (0.53)
+ Analysis (Ours) 3.72 (0.54) 4.73 (0.44)
Ablation Studies
We conduct ablation experiments to examine the contribution of each module in our multi-stage pipeline. Starting with a base that uses only the full paper text (Paper), we incrementally introduce components in the order they are executed: overall plan, architecture design, core logic, configuration file, and finally the analysis module. As shown in Table 6, performance steadily improves as additional steps are incorporated. Performance temporarily dips when the architecture design module is added. This can be attributed to the lack of dependency-aware ordering at this stage. For example, generating main.py before defining utils.py, which is only resolved later during the core logic phase. Once core logic is introduced, which defines file dependencies and generation order, scores improve substantially. Subsequent additions, such as the configuration file and final analysis stage, further refine the output. Our full system (Ours) achieves the highest scores under both reference-based (3.72) and reference-free (4.73) evaluation settings. These findings highlight the importance of sequential structuring in our pipeline.
Analysis on Executability
To verify whether the generated code is not only structurally sound but also executable with minimal intervention, we conduct a manual debugging analysis on five representative papers. For each case, we attempts to execute the generated repository and records the number of lines modified to achieve a successful run. Table 7 describes the total number of lines, the modified lines required, and the percentage of changes per repository. Remarkably, we find that on average, only 0.48% of the code needs to be altered for successful execution. Most modifications involve routine fixes such as updating deprecated OpenAI API calls to their latest versions or correcting simple type conversions. These findings indicate that our system does not merely produce syntactically correct or superficially plausible code, but outputs repositories that are functionally robust and executable with minimal human correction.
6 Conclusion and Limitations
6.1 Conclusion
In this work, we introduce PaperCoder, the framework designed to automatically generate code repositories from machine learning research papers. The framework follows a three-stage pipeline planning, analysis, and coding. The planning phase includes the construction of an overall plan, architectural design via diagrams, core logic planning with file dependencies, and configuration file generation. This plan is then refined through a detailed analysis of per-file implementation requirements, followed by sequential code generation guided by the planning outputs.
We evaluate PaperCoder on a comprehensive experimental Paper2Code benchmark consisting of 90 recent papers from top-tier machine learning conferences, as well as the PaperBench Code-Dev benchmark. Our framework consistently outperforms existing baselines across model-based evaluation metrics. In addition, we conduct human evaluations with graduate-level computer science researchers, who assess the quality of repositories generated for their own papers. Results show that 77% of participants preferred PaperCoder’s implementation over alternatives, and 83% found the outputs practically useful for real-world usage. These findings collectively demonstrate the effectiveness of PaperCoder in bridging the gap between research ideas and working code.
6.2 Limitations
While our PaperCoder shows strong performance on recent machine learning papers, its current scope is limited to this specific domain. Extending the framework to support a broader range of scientific fields remains an important direction for future work. Additionally, our primary evaluation relies on model-based metrics, which may not fully capture executability or runtime correctness. Although we perform case studies involving manual debugging to validate executability, a scalable and automated approach to execution-based evaluation—including fault localization and debugging—is a promising avenue for future research. Incorporating such capabilities would further enhance PaperCoder’s utility in real-world research workflows.
Appendix A Additional Experimental Designs
A.1 Implementation Details
All experiments are conducted using the o3-mini with high reasoning effort version (o3-mini-high) as the default backbone, released on January 31, 2025. To collect paper metadata and content, we utilize the openreview_scraper [^2] https://github.com/pranftw/openreview_scraper and the Semantic Scholar API [^3] https://www.semanticscholar.org/product/api. For document processing, we convert papers into structured JSON format using the s2orc-doc2json library [^4] https://github.com/allenai/s2orc-doc2json.
A.2 Human Evaluation Process
To conduct evaluations with human judges, we recruited 13 researchers from the South Korea, majoring in computer science, each with a minimum of 1 published papers. For annotation, they were provided with a 4-page document, which includes the task instructions, annotation examples, and 10 generated repositories grouped into three sets. Each repository was anonymized using a repoX naming format to prevent bias regarding the generation method. Following the question guidelines in the document, annotators reviewed and evaluated the repositories generated by different methods and models. Also, on average, evaluating 10 repositories for a single paper took approximately 45 minutes. Table 13 is a detailed annotation example.
A.3 Reference-Based Evaluation
In the reference-based evaluation setting, the repository may exceed the model’s context length. Following Starace et al. [40], when this occurs, we prompt the model to select the most relevant files for evaluation. The selected subset is then used as the reference for scoring. We use the gpt-4o-2024-11-20 as the evaluation model.
A.4 PaperBench Code-Dev Evaluation
While PaperCoder is designed to generate only the source code, the benchmark used for evaluation requires a reproduce.sh script. To meet this requirement, we additionally prompt the coding agent to generate a reproduce.sh file specifically for this benchmark, and we evaluate the repository including this file.
Table 7: Executability results across generated repositories. We manually select five papers and generate corresponding repositories using PaperCoder. For each repository, we report the number of lines modified during debugging, the total number of lines, and the percentage of modified lines.
Repo Name CoLoR cognitive-behaviors RADA Self-Instruct G-EVAL Average
Modified lines 2 1 1 25 10 6.5
Total lines 1132 2060 1609 1334 1374 1251.5
Percentage 0.18 0.05 0.06 1.87 0.73 0.48
Table 8: Qualitative analysis of top-ranked repositories. We categorize the reasons why human annotators selected the repositories generated by our framework as their top choice.
Completeness Clean Structure Faithfulness to Paper Ease of Use Code Quality Unique Strengths
8 6 5 4 2 3
Prompt for generating the overall plan in the planning stage [System] You are an expert researcher and strategic planner with a deep understanding of experimental design and reproducibility in scientific research. You will receive a research paper in JSON format. Your task is to create a detailed and efficient plan to reproduce the experiments and methodologies described in the paper. This plan should align precisely with the paper’s methodology, experimental setup, and evaluation metrics. Instructions: 1. Align with the Paper: Your plan must strictly follow the methods, datasets, model configurations, hyperparameters, and experimental setups described in the paper. 2. Be Clear and Structured: Present the plan in a well-organized and easy-to-follow format, breaking it down into actionable steps. 3. Prioritize Efficiency: Optimize the plan for clarity and practical implementation while ensuring fidelity to the original experiments. [User] ## Paper {paper_json} ## Task 1. We want to reproduce the method described in the attached paper. 2. The authors did not release any official code, so we have to plan our own implementation. 3. Before writing any Python code, please outline a comprehensive plan that covers: - Key details from the paper’s **Methodology**. - Important aspects of **Experiments**, including dataset requirements, experimental settings, hyperparameters, or evaluation metrics. 4. The plan should be as **detailed and informative** as possible to help us write the final code later. ## Requirements - You don’t need to provide the actual code yet; focus on a **thorough, clear strategy**. - If something is unclear from the paper, mention it explicitly. ## Instruction The response should give us a strong roadmap, making it easier to write the code later.
Figure 5: Prompt for generating the overall plan in the planning stage.
[…]
Figure 6: Prompt for generating the architecture design in the planning stage. This prompt follows the previous prompt and response shown in Figure 5.
Prompt for generating the logic design in the planning stage [User] Your goal is break down tasks according to PRD/technical design, generate a task list, and analyze task dependencies. You will break down tasks, analyze dependencies. You outline a clear PRD/technical design for reproducing the paper’s method and experiments. Now, let’s break down tasks according to PRD/technical design, generate a task list, and analyze task dependencies. The Logic Analysis should not only consider the dependencies between files but also provide detailed descriptions to assist in writing the code needed to reproduce the paper. —– ## Format Example [CONTENT] { "Required packages": [ "numpy==1.21.0", "torch==1.9.0" ], "Required Other language third-party packages": [ "No third-party dependencies required" ], "Logic Analysis": [ [ "data_preprocessing.py", "DataPreprocessing class …….." ], [ "trainer.py", "Trainer ……. " ], [ "dataset_loader.py", "Handles loading and …….." ], [ "model.py", "Defines the model ……." ], [ "evaluation.py", "Evaluation class …….. " ], [ "main.py", "Entry point ……." ] ], "Task list": [ "dataset_loader.py", "model.py", "trainer.py", "evaluation.py", "main.py" ], "Full API spec": "openapi: 3.0.0 …", "Shared Knowledge": "Both data_preprocessing.py and trainer.py share ……..", "Anything UNCLEAR": "Clarification needed on recommended hardware configuration for large-scale experiments." } [/CONTENT] ## Nodes: "<node>: <type> # <instruction>" - Required packages: typing.Optional[typing.List[str]] # Provide required third-party packages in requirements.txt format.(e.g., ’numpy==1.21.0’). - Required Other language third-party packages: typing.List[str] # List down packages required for non-Python languages. If none, specify "No third-party dependencies required". - Logic Analysis: typing.List[typing.List[str]] # Provide a list of files with the classes/methods/functions to be implemented, including dependency analysis and imports. Include as much detailed description as possible. - Task list: typing.List[str] # Break down the tasks into a list of filenames, prioritized based on dependency order. The task list must include the previously generated file list. - Full API spec: <class ’str’> # Describe all APIs using OpenAPI 3.0 spec that may be used by both frontend and backend. If front-end and back-end communication is not required, leave it blank. - Shared Knowledge: <class ’str’> # Detail any shared knowledge, like common utility functions or configuration variables. - Anything UNCLEAR: <class ’str’> # Mention any unresolved questions or clarifications needed from the paper or project scope. ## Constraint Format: output wrapped inside [CONTENT][/CONTENT] like the format example, nothing else. ## Action Follow the node instructions above, generate your output accordingly, and ensure it follows the given format example.
Figure 7: Prompt for generating the logic design in the planning stage. This prompt follows the previous prompt and response shown in Figure 6.
Prompt for generating the configuration file in the planning stage [User] You write elegant, modular, and maintainable code. Adhere to Google-style guidelines. Based on the paper, plan, design specified previously, follow the "Format Example" and generate the code. Extract the training details from the above paper (e.g., learning rate, batch size, epochs, etc.), follow the "Format example" and generate the code. DO NOT FABRICATE DETAILS — only use what the paper provides. You must write ‘config.yaml‘. ATTENTION: Use ’##’ to SPLIT SECTIONS, not ’#’. Your output format must follow the example below exactly. —– # Format Example ## Code: config.yaml “‘yaml ## config.yaml training: learning_rate: … batch_size: … epochs: … “‘ —– ## Code: config.yaml
Figure 8: Prompt for generating the configuration file in the planning stage. This prompt follows the previous prompt and response shown in Figure 7.
Prompt for analyzing [System] You are an expert researcher, strategic analyzer and software engineer with a deep understanding of experimental design and reproducibility in scientific research. You will receive a research paper in JSON format, an overview of the plan, a design in JSON format consisting of "Implementation approach", "File list", "Data structures and interfaces", and "Program call flow", followed by a task in JSON format that includes "Required packages", "Required other language third-party packages", "Logic Analysis", and "Task list", along with a configuration file named "config.yaml". Your task is to conduct a comprehensive logic analysis to accurately reproduce the experiments and methodologies described in the research paper. This analysis must align precisely with the paper’s methodology, experimental setup, and evaluation criteria. 1. Align with the Paper: Your analysis must strictly follow the methods, datasets, model configurations, hyperparameters, and experimental setups described in the paper. 2. Be Clear and Structured: Present your analysis in a logical, well-organized, and actionable format that is easy to follow and implement. 3. Prioritize Efficiency: Optimize the analysis for clarity and practical implementation while ensuring fidelity to the original experiments. 4. Follow design: YOU MUST FOLLOW "Data structures and interfaces". DONT CHANGE ANY DESIGN. Do not use public member functions that do not exist in your design. 5. REFER TO CONFIGURATION: Always reference settings from the config.yaml file. Do not invent or assume any values—only use configurations explicitly provided. [User]
# Context
## Paper {The content of the paper in json format} —– ## Overview of the plan {The content of the overall plan} —– ## Design {The content of the architecture design} —– ## Task {The content of the logic design} —– ## Configuration file “‘yaml {The content of the configuration file} “‘ —– ## Instruction Conduct a Logic Analysis to assist in writing the code, based on the paper, the plan, the design, the task and the previously specified configuration file (config.yaml). You DON’T need to provide the actual code yet; focus on a thorough, clear analysis. Write the logic analysis in ’{The name of the file to be generated"}’, which is intended for ’{Description of the file generated through the "Logic Analysis" step of the logic design.}’. —– ## Logic Analysis: {todo_file_name}
Figure 9: Prompts for analyzing. {} indicate placeholders to be filled with the content described in the accompanying explanation. The prompt is presented to the LLM for each file, following the sequence defined in the logic design.
Prompt for coding [System] You are an expert researcher and software engineer with a deep understanding of experimental design and reproducibility in scientific research. You will receive a research paper in JSON format, an overview of the plan, a Design in JSON format consisting of "Implementation approach", "File list", "Data structures and interfaces", and "Program call flow", followed by a Task in JSON format that includes "Required packages", "Required other language third-party packages", "Logic Analysis", and "Task list", along with a configuration file named "config.yaml". Your task is to write code to reproduce the experiments and methodologies described in the paper. The code you write must be elegant, modular, and maintainable, adhering to Google-style guidelines. The code must strictly align with the paper’s methodology, experimental setup, and evaluation metrics. Write code with triple quoto. [User]
# Context
## Paper {The content of the paper in json format} —– ## Overview of the plan {The content of the overall plan} —– ## Design {The content of the architecture design} —– ## Task {The content of the logic design} —– ## Configuration file “‘yaml {The content of the configuration file} “‘ —– ## Code Files {The content of the code files generated in the previous step.} —– # Format example ## Code: {todo_file_name} “‘python ## todo_file_name “‘ —– # Instruction Based on the paper, plan, design, task and configuration file(config.yaml) specified previously, follow "Format example", write the code. We have {done_file_lst}. Next, you must write only the "{todo_file_name}". 1. Only One file: do your best to implement THIS ONLY ONE FILE. 2. COMPLETE CODE: Your code will be part of the entire project, so please implement complete, reliable, reusable code snippets. 3. Set default value: If there is any setting, ALWAYS SET A DEFAULT VALUE, ALWAYS USE STRONG TYPE AND EXPLICIT VARIABLE. AVOID circular import. 4. Follow design: YOU MUST FOLLOW "Data structures and interfaces". DONT CHANGE ANY DESIGN. Do not use public member functions that do not exist in your design. 5. CAREFULLY CHECK THAT YOU DONT MISS ANY NECESSARY CLASS/FUNCTION IN THIS FILE. 6. Before using a external variable/module, make sure you import it first. 7. Write out EVERY CODE DETAIL, DON’T LEAVE TODO. 8. REFER TO CONFIGURATION: you must use configuration from "config.yaml". DO NOT FABRICATE any configuration values. {detailed_logic_analysis} ## Code: {todo_file_name}
Figure 10: Prompts for coding. {} indicate placeholders to be filled with the content described in the accompanying explanation. The prompt is presented to the LLM for each file, following the sequence defined in the logic design. Previously generated code files are accumulated and provided as part of the ## Code Files input.
Prompt for model-based reference-based evaluation [System] You will be given a research paper along with two corresponding code repositories: a gold repository and a target repository. Your task is to compare the target repository against the gold repository, rate the target repository on one metric, and provide a critique highlighting key differences. Please make sure you read and understand these instructions carefully. Keep this document open while reviewing, and refer to it as needed. Evaluation Criteria: Correctness (1-5): The quality of the target repository in accurately implementing the paper’s concepts, methodology, and algorithms without logical errors, as compared to the gold repository. Additionally, provide a critique focusing on the completeness, accuracy, and implementation choices made in the target repository relative to the gold repository. 1: Very Poor. The target repository does not correctly implement the core concepts, methodology, or algorithms from the paper. Major logical errors or missing components are present, especially when compared to the gold repository. 2: Poor. The target repository attempts to implement the paper’s concepts but contains significant mistakes or missing components, making the implementation incorrect when compared to the gold repository. 3: Fair. Some core components and concepts are correctly implemented in the target repository, but there are notable logical errors or inaccuracies compared to the gold repository. 4: Good. The target repository correctly implements the key components and methodology, with only minor inaccuracies or deviations from the gold repository. 5: Excellent. The target repository fully and accurately implements all relevant key components, methodology, and algorithms from the paper, matching the quality of the gold repository. Evaluation Steps 1. Identify Key Aspects of the Paper: Carefully read the research paper to understand its core concepts, methodology, and algorithms. Pay close attention to the key aspects that are crucial for implementing the paper’s results (e.g., specific algorithms, data preprocessing steps, evaluation protocols). 2. Analyze the Gold Repository: Examine the gold repository to understand how these key aspects have been implemented. Use the gold repository as a reference for how the paper’s methodology should be translated into code. Note the completeness, accuracy, and design choices in the gold repository that faithfully represent the paper’s concepts. 3. Examine the Target Repository: Analyze the target repository to assess how well it implements the key aspects of the paper. Reference the gold repository as a guide for understanding these key aspects in the target repository. Focus on whether the target repository’s core logic, algorithms, and structure align with the methodology and experiments described in the paper. 4. Identify Logical Errors and Deviations: Check for logical errors, missing steps, or deviations from the paper’s methodology. Note any incorrect representations, inconsistencies, or incomplete implementations that could affect the correctness of the target repository. 5. Provide a Critique: Consider both the completeness and accuracy of the implementation relative to the paper’s goals and the gold repository’s standard. You do not need to analyze minor details like logging functions, script organization, or documentation quality. Instead, concentrate on the correctness of the logic and implementation that ensures the core concepts from the paper are fully reflected in the target repository. For each mismatch or deviation in implementation, note down specific critiques comparing relevant functions in the target repository to the corresponding functions in the gold repository. Highlight incorrect logic, missing steps, or deviations that affect the correct implementation of the paper’s methodology. 5. Assess the Correctness: Determine whether the target repository includes all the critical elements described in the paper and implemented in the gold repository. Identify missing components, significant deviations, or incorrect implementations that could affect the correctness of the target repository. 6. Assign a Score: Based on your evaluation, provide a critique and assign a correctness score from 1 to 5 for the target repository, reflecting how well it implements the key aspects of the paper refer to the gold repository. Include a detailed critique in the specified JSON format. Severity Level: Each identified critique will be assigned a severity level based on its impact on the correctness of the methodology implementation. - High: Missing or incorrect implementation of the paper’s core concepts, major loss functions, or experiment components that are fundamental to reproducing the paper’s methodology. - Example: The main algorithm is missing or fundamentally incorrect. - Medium: Issues affecting training logic, data preprocessing, or other core functionalities that significantly impact performance but do not completely break the system. - Example: Improper training loop structure, incorrect data augmentation, or missing essential components in data processing. - Low: Errors in specific features that cause deviations from expected results but can be worked around with modifications. Any errors in the evaluation process belong to this category unless they impact the core concepts. These include minor issues like logging, error handling mechanisms, configuration settings, evaluation steps that do not alter the fundamental implementation and additional implementations not explicitly stated in the paper. - Example: Suboptimal hyperparameter initialization, incorrect learning rate schedule, inaccuracies in evaluation metrics, using a different random seed, variations in batch processing, different weight initialization, issues in result logging or reporting, variations in evaluation dataset splits, improper error handling in non-critical steps, mismatches in secondary evaluation criteria, or additional implementation details not specified in the paper that do not interfere with core results.
Prompt for model-based reference-based evaluation Example JSON format: “‘json { "critique_list": [ { "gold_file_name": "preprocessing.py", "gold_func_name": "data_process", "target_file_name": "dataset.py", "target_func_name": "train_preprocess", "severity_level": "medium", "critique": "A critique of the target repository’s file with reference to the gold repository." }, { "gold_file_name": "utils.py", "gold_func_name": "calculate_metric", "target_file_name": "metric.py", "target_func_name": "f1_at_k" "severity_level": "low", "critique": "A critique of the target repository’s file with reference to the gold repository." }, ], "score": 2 } “‘ Sample: Research Paper: {{The content of the paper}} Gold Repository: {{The gold repository, officially released by the authors, serves as the reference implementation.}} Target Repository: {{The generated repository, which serves as the target repository for evaluation.}} Please provide critique of the target repository and a single numerical rating (1, 2, 3, 4, or 5) based on the quality of the sample, following the Example JSON format, without any additional commentary, formatting, or chattiness.
Figure 11: Prompts for model-based reference-based evaluation. {{}} indicate placeholders to be filled with the content described in the accompanying explanation.
Prompt for model-based reference-free evaluation [System] You will be given a research paper along with its corresponding code repository. Your task is to rate the code repository on one metric and provide a critique highlighting key differences. Please make sure you read and understand these instructions carefully. Keep this document open while reviewing, and refer to it as needed. Evaluation Criteria: Correctness (1-5): The quality of the repository in accurately implementing the paper’s concepts, methodology, and algorithms without logical errors. Additionally, provide a critique focusing on the completeness, accuracy, and implementation choices made in the repository relative to the methodology and algorithms described in the paper. 1: Very Poor. The repository does not correctly implement the core concepts, methodology, or algorithms from the paper. Major logical errors or missing components are present. 2: Poor. The repository attempts to implement the paper’s concepts but contains significant mistakes or missing components, making the implementation incorrect. 3: Fair. Some core components and concepts are correctly implemented, but there are notable logical errors or inaccuracies in the methodology. 4: Good. The repository correctly implements the key components and methodology, with only minor inaccuracies that do not significantly affect correctness. 5: Excellent. The repository fully and accurately implements all key components, methodology, and algorithms from the paper without logical errors. Evaluation Steps 1. Identify Key Aspects of the Paper: Carefully read the paper to understand its core concepts, methodology, and algorithms. Pay close attention to key aspects crucial for implementing the paper’s results (e.g., specific algorithms, data preprocessing steps, evaluation protocols). 2. Examine the Code Repository: Analyze the repository to determine how well it implements the key aspects of the paper. Focus on whether the repository’s core logic, algorithms, and structure align with the methodology and experiments described in the paper. 3. Identify Logical Errors and Deviations: Check for logical errors, missing steps, or deviations from the paper’s methodology. Note any incorrect representations, inconsistencies, or incomplete implementations that could affect the correctness of the repository. 4. Provide a Critique: Consider the completeness and accuracy of the implementation relative to the paper’s goals. You do not need to analyze minor details like logging functions, script organization, or documentation quality. Instead, concentrate on the correctness of the logic and implementation to ensure the core concepts from the paper are fully reflected in the repository. For each identified issue, write a detailed critique specifying the affected files and functions in the repository. Highlight missing or incorrectly implemented steps that impact the correctness and alignment with the paper’s methodology. 5. Assess Completeness and Accuracy: Evaluate the repository for its completeness and accuracy relative to the paper’s methodology. Ensure that all critical components—such as data preprocessing, core algorithms, and evaluation steps—are implemented and consistent with the paper’s descriptions. 6. Assign a Score: Based on your evaluation, provide a critique and assign a correctness score from 1 to 5 for the repository, reflecting how well it implements the key aspects of the paper. Include a detailed critique in the specified JSON format. Severity Level: Each identified critique will be assigned a severity level based on its impact on the correctness of the methodology implementation. - High: Missing or incorrect implementation of the paper’s core concepts, major loss functions, or experiment components that are fundamental to reproducing the paper’s methodology. - Example: The main algorithm is missing or fundamentally incorrect. - Medium: Issues affecting training logic, data preprocessing, or other core functionalities that significantly impact performance but do not completely break the system. - Example: Improper training loop structure, incorrect data augmentation, or missing essential components in data processing. - Low: Errors in specific features that cause deviations from expected results but can be worked around with modifications. Any errors in the evaluation process belong to this category unless they impact the core concepts. These include minor issues like logging, error handling mechanisms, configuration settings, evaluation steps that do not alter the fundamental implementation and additional implementations not explicitly stated in the paper. - Example: Suboptimal hyperparameter initialization, incorrect learning rate schedule, inaccuracies in evaluation metrics, using a different random seed, variations in batch processing, different weight initialization, issues in result logging or reporting, variations in evaluation dataset splits, improper error handling in non-critical steps, mismatches in secondary evaluation criteria, or additional implementation details not specified in the paper that do not interfere with core results.
Prompt for model-based reference-free evaluation Example JSON format: “‘json { "critique_list": [ { "file_name": "dataset.py", "func_name": "train_preprocess", "severity_level": "medium", "critique": "A critique of the target repository’s file." }, { "file_name": "metrics.py", "func_name": "f1_at_k", "severity_level": "low", "critique": "A critique of the target repository’s file." } ], "score": 2 } “‘ Sample: Research Paper: {{The content of the paper}} Code Repository: {{The generated repository, which serves as the target repository for evaluation.}} Please provide a critique list for the code repository and a single numerical rating (1, 2, 3, 4, or 5) based on the quality of the sample, following the Example JSON format, without any additional commentary, formatting, or chattiness.
Figure 12: Prompts for model-based reference-free evaluation. {{}} indicate placeholders to be filled with the content described in the accompanying explanation.
Table 9: List of ICLR 2024 papers used in our experimental benchmark. We evaluate each paper using the model-based, reference-free setting, with gpt-4o-2024-11-20 as the evaluation model.
Paper
Source Score
Generative Judge for Evaluating Alignment
Poster 4
Distributional Preference Learning: Understanding and Accounting for Hidden Context in RLHF
Poster 4
Inherently Interpretable Time Series Classification via Multiple Instance Learning
Oral 3.9
iTransformer: Inverted Transformers Are Effective for Time Series Forecasting
Oral 3.9
Tell Your Model Where to Attend: Post-hoc Attention Steering for LLMs
Poster 3.9
Knowledge Distillation Based on Transformed Teacher Matching
Poster 3.9
Meaning Representations from Trajectories in Autoregressive Models
Poster 3.8
A Simple Interpretable Transformer for Fine-Grained Image Classification and Analysis
Poster 3.8
VDC: Versatile Data Cleanser based on Visual-Linguistic Inconsistency by Multimodal Large Language Models
Poster 3.8
Vocos: Closing the gap between time-domain and Fourier-based neural vocoders for high-quality audio synthesis
Poster 3.8
SliceGPT: Compress Large Language Models by Deleting Rows and Columns
Poster 3.8
Beyond Accuracy: Evaluating Self-Consistency of Code Large Language Models with IdentityChain
Poster 3.8
Guiding Masked Representation Learning to Capture Spatio-Temporal Relationship of Electrocardiogram
Poster 3.8
Social Reward: Evaluating and Enhancing Generative AI through Million-User Feedback from an Online Creative Community
Oral 3.7
Language Model Detectors Are Easily Optimized Against
Poster 3.7
Improving protein optimization with smoothed fitness landscapes
Poster 3.7
SparseFormer: Sparse Visual Recognition via Limited Latent Tokens
Poster 3.7
AutoVP: An Automated Visual Prompting Framework and Benchmark
Poster 3.7
Hierarchical Context Merging: Better Long Context Understanding for Pre-trained LLMs
Poster 3.7
SEABO: A Simple Search-Based Method for Offline Imitation Learning
Poster 3.7
OpenChat: Advancing Open-source Language Models with Mixed-Quality Data
Poster 3.7
Rethinking The Uniformity Metric in Self-Supervised Learning
Poster 3.7
VONet: Unsupervised Video Object Learning With Parallel U-Net Attention and Object-wise Sequential VAE
Poster 3.6
Efficient Backpropagation with Variance-Controlled Adaptive Sampling
Poster 3.6
Structuring Representation Geometry with Rotationally Equivariant Contrastive Learning
Poster 3.6
ControlVideo: Training-free Controllable Text-to-Video Generation
Poster 3.6
Context-Aware Meta-Learning
Poster 3.6
RECOMBINER: Robust and Enhanced Compression with Bayesian Implicit Neural Representations
Poster 3.6
Peering Through Preferences: Unraveling Feedback Acquisition for Aligning Large Language Models
Poster 3.6
Modulate Your Spectrum in Self-Supervised Learning
Poster 3.6
Table 10: List of NeurIPS 2024 papers used in our experimental benchmark. We evaluate each paper using the model-based, reference-free setting, with gpt-4o-2024-11-20 as the evaluation model.
Paper
Source Score
PACE: marrying generalization in PArameter-efficient fine-tuning with Consistency rEgularization
Oral 4
The Road Less Scheduled
Oral 4
G-Retriever: Retrieval-Augmented Generation for Textual Graph Understanding and Question Answering
Poster 4
Binarized Diffusion Model for Image Super-Resolution
Poster 4
Learning to Predict Structural Vibrations
Poster 4
Attack-Aware Noise Calibration for Differential Privacy
Poster 4
Make Your LLM Fully Utilize the Context
Poster 3.9
Smoothed Energy Guidance: Guiding Diffusion Models with Reduced Energy Curvature of Attention
Poster 3.9
Sm: enhanced localization in Multiple Instance Learning for medical imaging classification
Poster 3.9
AutoTimes: Autoregressive Time Series Forecasters via Large Language Models
Poster 3.9
End-to-End Ontology Learning with Large Language Models
Poster 3.8
Scaling transformer neural networks for skillful and reliable medium-range weather forecasting
Poster 3.8
Autoregressive Image Generation without Vector Quantization
Oral 3.7
Adaptive Randomized Smoothing: Certified Adversarial Robustness for Multi-Step Defences
Oral 3.7
Generalizable Person Re-identification via Balancing Alignment and Uniformity
Poster 3.7
Universal Neural Functionals
Poster 3.7
Are Self-Attentions Effective for Time Series Forecasting?
Poster 3.7
xMIL: Insightful Explanations for Multiple Instance Learning in Histopathology
Poster 3.7
Leveraging Environment Interaction for Automated PDDL Translation and Planning with Large Language Models
Poster 3.7
Task-Agnostic Machine Learning-Assisted Inference
Poster 3.7
Make Continual Learning Stronger via C-Flat
Poster 3.7
DARG: Dynamic Evaluation of Large Language Models via Adaptive Reasoning Graph
Poster 3.7
AsyncDiff: Parallelizing Diffusion Models by Asynchronous Denoising
Poster 3.7
You Only Look Around: Learning Illumination Invariant Feature for Low-light Object Detection
Poster 3.6
MutaPLM: Protein Language Modeling for Mutation Explanation and Engineering
Poster 3.6
Advancing Training Efficiency of Deep Spiking Neural Networks through Rate-based Backpropagation
Poster 3.6
Improved off-policy training of diffusion samplers
Poster 3.6
Navigating the Effect of Parametrization for Dimensionality Reduction
Poster 3.6
Long-Range Feedback Spiking Network Captures Dynamic and Static Representations of the Visual Cortex under Movie Stimuli
Poster 3.6
InfLLM: Training-Free Long-Context Extrapolation for LLMs with an Efficient Context Memory
Poster 3.6
Table 11: List of ICML 2024 papers used in our experimental benchmark. We evaluate each paper using the model-based, reference-free setting, with gpt-4o-2024-11-20 as the evaluation model.
Paper
Source Score
SAMformer: Unlocking the Potential of Transformers in Time Series Forecasting with Sharpness-Aware Minimization and Channel-Wise Attention
Oral 4
Autoformalizing Euclidean Geometry
Poster 4
Recurrent Distance Filtering for Graph Representation Learning
Poster 4
CosPGD: an efficient white-box adversarial attack for pixel-wise prediction tasks
Poster 3.9
Token-level Direct Preference Optimization
Poster 3.9
BayOTIDE: Bayesian Online Multivariate Time Series Imputation with Functional Decomposition
Oral 3.8
CurBench: Curriculum Learning Benchmark
Poster 3.8
Exploring the Low-Pass Filtering Behavior in Image Super-Resolution
Poster 3.8
Towards Efficient Exact Optimization of Language Model Alignment
Poster 3.7
On the Effectiveness of Supervision in Asymmetric Non-Contrastive Learning
Poster 3.7
Drug Discovery with Dynamic Goal-aware Fragments
Poster 3.7
Fool Your (Vision and) Language Model With Embarrassingly Simple Permutations
Poster 3.7
Image Restoration Through Generalized Ornstein-Uhlenbeck Bridge
Poster 3.7
Timer: Generative Pre-trained Transformers Are Large Time Series Models
Poster 3.7
Mitigating Oversmoothing Through Reverse Process of GNNs for Heterophilic Graphs
Poster 3.7
Scribble-Supervised Semantic Segmentation with Prototype-based Feature Augmentation
Poster 3.7
ConvNet vs Transformer, Supervised vs CLIP: Beyond ImageNet Accuracy
Poster 3.7
CLIF: Complementary Leaky Integrate-and-Fire Neuron for Spiking Neural Networks
Oral 3.6
FiT: Flexible Vision Transformer for Diffusion Model
Oral 3.6
Decomposing Uncertainty for Large Language Models through Input Clarification Ensembling
Oral 3.6
SparseTSF: Modeling Long-term Time Series Forecasting with *1k* Parameters
Oral 3.6
Sample-specific Masks for Visual Reprogramming-based Prompting
Oral 3.6
Boundary Exploration for Bayesian Optimization With Unknown Physical Constraints
Poster 3.6
Listwise Reward Estimation for Offline Preference-based Reinforcement Learning
Poster 3.6
Graph Distillation with Eigenbasis Matching
Poster 3.6
Temporal Spiking Neural Networks with Synaptic Delay for Graph Reasoning
Poster 3.6
Position: Quo Vadis, Unsupervised Time Series Anomaly Detection?
Poster 3.6
Neural SPH: Improved Neural Modeling of Lagrangian Fluid Dynamics
Poster 3.6
Self-Play Fine-Tuning Converts Weak Language Models to Strong Language Models
Poster 3.6
Unveiling and Harnessing Hidden Attention Sinks: Enhancing Large Language Models without Training through Attention Calibration
Poster 3.6
Table 12: List of papers used in human evaluation. We evaluate each paper using the model-based, reference-free setting, with gpt-4o-2024-11-20 as the evaluation model.
Paper
Score
VideoICL: Confidence-based Iterative In-context Learning for Out-of-Distribution Video Understanding
2.6
Identity Decoupling for Multi-Subject Personalization of Text-to-Image Models
3.3
Knowledge-Augmented Language Model Verification
3.3
SEA: Sparse Linear Attention with Estimated Attention Mask
2.8
HarmAug: Effective Data Augmentation for Knowledge Distillation of Safety Guard Models
3.0
Graph Generation with Diffusion Mixture
3.7
Adaptive-RAG: Learning to Adapt Retrieval-Augmented Large Language Models through Question Complexity
2.9
Sketch-of-Thought: Efficient LLM Reasoning with Adaptive Cognitive-Inspired Sketching
4.0
Mol-LLaMA: Towards General Understanding of Molecules in Large Molecular Language Model
3.5
Rethinking Code Refinement: Learning to Judge Code Efficiency
3.1
Knowledge-Augmented Reasoning Distillation for Small Language Models in Knowledge-Intensive Tasks
3.2
Concept-skill Transferability-based Data Selection for Large Vision-Language Models
3.0
Aligning to thousands of preferences via system message generalization
3.5
Table 13: List of papers used in executability analysis.
Repo Name
Paper
CoLoR
Efficient Long Context Language Model Retrieval with Compression
cognitive-behaviors
Cognitive Behaviors that Enable Self-Improving Reasoners, or, Four Habits of Highly Effective STaRs
RADA
Retrieval-Augmented Data Augmentation for Low-Resource Domain Tasks
Self-Instruct
Self-Instruct: Aligning Language Models with Self-Generated Instructions
G-EVAL
G-Eval: NLG Evaluation using GPT-4 with Better Human Alignment
Figure 13: Human Evaluation Format
Figure 14: Human Evaluation Format
Figure 15: Human Evaluation Format