### v3

# Paper2Code: Automating Code Generation from Scientific Papers in Machine Learning

Abstract
Despite the rapid growth of machine learning research, corresponding code implementations are often unavailable, making it slow and labor-intensive for researchers to reproduce results and build upon prior work. In the meantime, recent Large Language Models (LLMs) excel at understanding scientific documents and generating high-quality code. Inspired by this, we introduce PaperCoder, a multi-agent LLM framework that transforms machine learning papers into functional code repositories. PaperCoder operates in three stages: planning, where it constructs a high-level roadmap, designs the system architecture with diagrams, identifies file dependencies, and generates configuration files; analysis, which focuses on interpreting implementation-specific details; and generation, where modular, dependency-aware code is produced. Moreover, each phase is instantiated through a set of specialized agents designed to collaborate effectively across the pipeline. We then evaluate PaperCoder on generating code implementations from machine learning papers based on both model-based and human evaluations, particularly from the authors of those papers, with author-released repositories as ground truth if available. Our results demonstrate the effectiveness of PaperCoder in creating high-quality, faithful implementations. Furthermore, it consistently shows strengths in the recently released PaperBench benchmark, surpassing strong baselines by substantial margins. Code is available at: https://github.com/going-doer/Paper2Code.
1 Introduction
Reproducibility lies at the heart of scientific progress, which enables researchers to validate findings, build upon prior work, and ultimately push the boundaries of knowledge [6, 3, 30]. However, across many disciplines, reproducing scientific results remains an enduring challenge. This is often due to incomplete documentation, missing experimental details, lack of access to data or proprietary tools, and, especially in machine learning research, the absence of corresponding code: for example, only average 19.5% of the papers accepted to top-tier machine learning conferences in 2024 provide their code implementations shown in Figure 1. As a result, researchers frequently invest substantial effort in reverse-engineering methods and experimental results from papers, a process that is both time-consuming and labor-intensive, subsequently slowing down the overall pace of scientific innovation.
Meanwhile, recent Large Language Models (LLMs) have shown outstanding capabilities in understanding and generating both natural language and programming code [10, 28, 35], with performances increasingly approaching or even surpassing that of domain experts in some scenarios. In addition, this progress has sparked growing interest in leveraging LLMs to accelerate scientific workflows, particularly in the early stages of ideation for new and valid research hypotheses [24, 47, 36, 2, 18, 48, 37]. Furthermore, some of these studies, as well as others focusing on later stages of automating experimental validations and improvements [14, 51, 4, 41], demonstrate the potential of LLMs to generate code and even carry out experiments end-to-end; however, they typically assume and heavily rely on access to pre-existing implementations, partial code snippets, or well-defined APIs. As such, it remains questionable whether generating complete and faithful implementations solely from research papers (without access to prior code, APIs, or additional supplementary materials) can be achievable.
To answer this question, we introduce PaperCoder, a multi-agent LLM-powered framework, designed to automatically generate faithful code repositories in machine learning directly from and contextualized with research papers, which differs from prior work that requires partial implementations from human inputs. Specifically, PaperCoder aims to emulate the typical life cycle of human developers and researchers in writing the repository-level code, by decomposing the task into three structured stages: planning, analysis, and generation. First, during the planning stage, the proposed framework constructs a high-level roadmap to identify core components to implement, draws the overall system architecture with class and sequence diagrams to model structural relationships between modules, identifies file dependencies with their execution orders to guide correct build and execution flows, and generates configuration files to enable flexible customization of experimental workflows by human researchers. This is followed by the analysis stage, performing a fine-grained interpretation of each file and function with respect to their intended functionality, such as required inputs and outputs, interactions with other modules, and any algorithmic or architectural constraints derived from the source paper. Finally, in the generation stage, the framework synthesizes the entire code base based on the execution order determined earlier, along with the artifacts produced in the previous stages.
(a) PaperCoder overview
(b) Code availability
Figure 1: (a) PaperCoder overview. The proposed PaperCoder aims to transform given scientific papers (in machine learning domains) into code repositories, which consists of three sequential steps: planning, analysis, and coding. (b) Code availability, where blue bars indicate the total number of accepted papers and orange regions show those with officially released code (See Appendix B.1 for details on the availability calculation).
To validate the effectiveness of PaperCoder, we conduct extensive evaluations on a subset of recent machine learning papers accepted at top-tier venues in 2024, including ICLR, ICML, and NeurIPS referred to as our proposed Paper2Code benchmark (in short, Paper2CodeBench). Also, we incorporate the recent PaperBench benchmark [38] in our evaluation suite, enabling fine-grained evaluations of code implementations with some papers from ICML 2024. Then, on a battery of tests conducted not only with automated model-based evaluations (covering both reference-free and reference-based settings, conditional on the availability of author-released ground-truth repositories) but also with expert human evaluations (based on authors of original papers), PaperCoder demonstrates substantial improvements over baselines, generating more valid and faithful code repositories that could meaningfully support human researchers in understanding and reproducing prior work. Specifically, 88% of the generated repositories by PaperCoder are rated as the best over baselines, and 92% of human judges report that the generated repositories are indeed helpful. Also, further analyses show that each component of PaperCoder (consisting of planning, analysis, and generation) contributes to the performance gains, but also that the generated codebases can be executed, sometimes with only minor modifications (averaging 0.81% of total code lines) in cases where execution errors occur.
2 Related Work
Large Language Models for Code
LLMs have shown impressive capabilities in text understanding and generation [28, 10, 35], which have been widely utilized within specialized domains (beyond general tasks), such as mathematics, science, and coding [32, 43, 40]. Particularly, code-specialized LLMs [9, 8, 15] have received significant attention thanks to remarkable performance on various software engineering tasks [45], including software design and development [34, 13], requirements elicitation [27], and formal specification generation [25]. Our work aligns closely with this line of research, exploring and expanding upon the capabilities and applications of (code-specialized) LLMs.
Repository-Level Coding
Early work on LLM-based code generation typically focuses on single-file tasks, whose objective is to generate relatively short code snippets to solve isolated tasks, such as (algorithmic-level) programming competition problems or toy-level programming queries [19, 5, 1, 12]. However, as LLMs have advanced in comprehending and generating code with the ability to perform long-context reasoning, recent studies have increasingly shifted their attention toward more challenging repository-level coding tasks, which involve generating multi-file repositories that jointly account for architectural design, modular structure, and inter-file dependencies [21, 16, 39]. In particular, several recent efforts explore this emerging paradigm [50, 29], often adopting multi-agent or role-based frameworks to emulate realistic development workflows. For instance, ChatDev instantiates LLMs into role-playing agents that collaborate through structured dialogues [34], while MetaGPT implements a waterfall-style development pipeline with specialized agents for handling different stages of software engineering [13]. Beyond prior work, we explore the underexplored yet challenging task of transforming full-length, complex papers into repository-level implementations.
LLM-Powered Scientific Research
LLMs have been increasingly adopted to support every stage of the scientific process from ideation to experimental validation [31, 2, 18, 48, 33, 7, 20, 44]; thereby, helping researchers overcome existing challenges and ultimately accelerate scientific discovery [17, 24, 47]. Specifically, in fields such as computer science and machine learning (where code-based experimentation is central), LLMs have been used to design, refine, and extend code implementations. Yet, many recent efforts in this space assume access to and build on top of the original codebase [14, 46, 4, 41], which significantly limits their applicability in real-world scenarios since such implementations are oftentimes unavailable (Figure 1). To address this, concurrent to our work, Starace et al. [38] introduces a benchmark dataset called PaperBench, evaluating the capability of existing agentic AI systems in reproducing papers with fine-grained metrics. Notably, on top of PaperBench (which emphasizes evaluation), we further complement and extend this line by focusing on methodological aspects of how to transform scientific papers into repository-level code implementations.
3 Method
In this section, we start with describing the task of repository-level code generation from machine learning papers, and propose PaperCoder, a multi-agent, multi-stage framework designed to tackle it.
3.1 Repository-Level Code Generation from Scientific Papers in Machine Learning
The goal of our repository-level code generation task is to automatically produce a code repository that faithfully implements the methods and experiments described in machine learning papers (especially for cases where authors do not release their code), to support reproducibility and accelerate scientific progress [26, 30]. Formally, we define this task as a function (or a model) $M$ that maps a research paper $R$ to a corresponding code repository $C$, as follows: $M(R)=C$. Here, $C$ is composed of multiple interdependent files $\{c_{1},c_{2},...,c_{n}\}$, each responsible for implementing different components of the methods and experiments described in $R$, but together they should form a cohesive pipeline.
It is worth noting that the most straightforward approach to instantiating $M$ is to instruct the LLM to generate the entire code repository, conditioned on the given paper, as follows: $M(R):=\texttt{LLM}(\mathcal{T}(R))$, where $\mathcal{T}$ is the prompt template that specifies the intended behavior of the LLM for the target task (including task descriptions, detailed instructions, and any other relevant context). Yet, generating a complete, modular, and faithful repository (spanning multiple files) in a single pass is extremely challenging, even for powerful LLMs, due to the inherent complexity of scientific papers and their corresponding implementations, the long-context limitations of current models, and the difficulty in maintaining consistent global structure and cross-file dependencies. Therefore, we propose a multi-stage and multi-agent framework that decomposes the overall task into smaller, manageable subtasks, each handled by a specialized agent tailored to a specific aspect of paper-to-code transformation.
Figure 2: (Left) The naive approach, which directly generates an entire code repository from a paper. (Right) Our PaperCoder framework, which is operationalized by decomposing the task into three stages: (1) Planning, where a high-level implementation plan is constructed from the paper, including overall plan, architectural design, logic design, and configuration file; (2) Analysis, where the plan is translated into detailed file-level specifications; and (3) Coding, where the final codes are generated to implement the methods and experiments of the paper.
3.2 PaperCoder: LLM-Powered Multi-Agent Framework for Paper-to-Code Transformation
We now introduce PaperCoder, a structured, multi-agent framework for generating code repositories directly from machine learning papers (without access to pre-existing artifacts or implementations, such as skeleton code). Specifically, inspired by typical software development workflows, PaperCoder decomposes the task into three coordinated stages: Planning, Analysis, and Coding, each orchestrated by specialized LLM agents. Formally, given a paper $R$, the overall process can be defined as follows:
$\text{{Planning: }}P=M_{\text{plan}}(R),\quad\text{{Analysis: }}A=M_{\text{analysis}}(R,P),\quad\text{{Coding: }}C=M_{\text{code}}(R,P,A),$
where $P$, AA, and $C$ represent the high-level implementation plan, the detailed function-level analysis, and the final code repository, respectively. The overall pipeline of PaperCoder is shown in Figure 2.
3.2.1 Planning
It is worth noting that, in contrast to implementation specifications designed explicitly for software development, scientific papers are written to communicate ideas and findings to humans. As a result, they often contain high-level motivations, persuasive narratives, and auxiliary details that are crucial for human understanding but noisy, loosely specified, or ambiguous from a software engineering perspective. To mitigate this challenge, we introduce a structured planning phase that transforms unstructured textual content into implementation-level abstractions. Also, we decompose the planning process into four sequential subcomponents (to simplify the task and reduce cognitive and computational load of LLM-powered agents at each step): 1) overall plan, 2) architecture design, 3) logic design, and 4) configuration generation. Formally, we define this as: $M_{\text{plan}}(R)\rightarrow P=\{o,d,l,g\}$, where $o$ is the overall plan, $d$ is the architecture design, $l$ is the logic design, and $g$ is the configuration file, with each stage using the outputs of the previous ones as contextual input. We then describe how each subcomponent is instantiated through specific objectives and instructions below.
Overall Plan
The first step is to extract a high-level summary of the core components and functionalities described throughout the paper, to identify the specific methods and experiments to be implemented. In other words, this high-level overview includes model components, training objectives, data processing steps, and evaluation protocols (distributed across the entire paper), which can form the foundation for all subsequent steps, formalized as follows: $M_{\text{plan}}^{(1)}(R):=\texttt{LLM}(\mathcal{T}_{\text{plan}}^{(1)}(R))\rightarrow o$.
Architecture Design
Based on the extracted overall plan alongside the input paper, the next step is to define the repository-level architecture, which includes identifying files, organizing them into modules, and defining their relationships, to ensure a coherent and maintainable structure. Specifically, the LLM-powered agent is prompted to generate a file list, which outlines the overall file structure of the repository; a class diagram, which details static representations of files (such as core classes and their attributes); and a sequence diagram, which models the dynamic interactions. Formally, similar to overall plan, this process can be defined as follows: $M_{\text{plan}}^{(2)}(R,o):=\texttt{LLM}(\mathcal{T}_{\text{plan}}^{(2)}(R,o))\rightarrow d$.
Logic Design
While the previous architecture design focuses on what to build, the logic design phase specifies how these components should be instantiated in practice by considering their dependencies in terms of overall execution flow. In other words, its primary goal is to translate abstract component-level relationships from the architecture design into an actionable execution order for the entire implementation. This step is crucial because individual modules often depend on shared utilities, configurations, or data loaders that are defined in other parts of the repository, and without an explicitly defined execution order, the code generation can result in failure or inconsistency (e.g., generating file B before file A when B imports modules from A). To address this, the logic design stage not only produces an ordered file list that dictates the sequence in which the files should be implemented and executed, but also further elaborates on the logic within each file; thereby, providing more fine-grained specifications. Formally, $M_{\text{plan}}^{(3)}(R,o,d):=\texttt{LLM}(\mathcal{T}_{\text{plan}}^{(3)}(R,o,d))\rightarrow l$.
Configuration Generation
In the last stage of planning, PaperCoder synthesizes a configuration file (namely, config.yaml) that includes key hyperparameters, model settings, and other runtime options based on prior outputs alongside the given paper. We note that, in addition to grounding the code generation process with the explicit configuration details, it enables researchers to easily review and adjust experimental configurations without modifying the source code. Formally, $M_{\text{plan}}^{(4)}(R,o,d,l):=\texttt{LLM}(\mathcal{T}_{\text{plan}}^{(4)}(R,o,d,l))\rightarrow g$. We provide prompts used to elicit each planning output in Appendix E.
3.2.2 Analysis
Following the planning stage, which defines the overall structure and execution flow of the repository, the analysis phase focuses on interpreting and specifying the implementation-level details for modules within each file. In other words, unlike planning that answers what components to build and how they relate, this phase addresses the question of how each component should be operationalized and concretely implemented at the file level, which includes the definition of functional goals, input-output behaviors, intra- and inter-file dependencies, and algorithmic specifications derived from the original paper. Specifically, given the input paper $R$ and planning outputs $P=\{o,d,l,g\}$, the analysis agent iteratively processes each file $f_{i}$ (identified during planning) and generates a detailed analysis $a_{i}$ describing what needs to be implemented in that file. Formally, $\{M_{\text{analysis}}(R,P,f_{i})\}_{i=1}^{n=|F|}$ where $M_{\text{analysis}}(R,P,f_{i}):=\texttt{LLM}(\mathcal{T}_{\text{analysis}}(R,P,f_{i}))\rightarrow a_{i}$, with $F$ as the set of identified files, e.g., $f_{i}\in F$.
3.2.3 Coding
The final stage of our PaperCoder framework is the coding phase, where the complete code repository is produced. In particular, each file is generated based on all the available contextual information accumulated from the previous stages, including the overall plan, architecture design, logic design, configuration file, and file-specific analyses, as well as the original paper. Additionally, to ensure consistency across different files, we generate them sequentially according to the execution order (i.e., the ordered file list determined during the logic design stage). To be formal, the overall code generation process is defined as follows: for each file $f_{i}$, the corresponding code $c_{i}$ is generated as $M_{\text{code}}(R,P,f_{i},a_{i},\{c_{1},...,c_{i-1}\}):=\texttt{LLM}(\mathcal{T}_{\text{code}}(R,P,f_{i},a_{i},\{c_{1},...,c_{i-1}\}))\rightarrow c_{i}$, resulting in the complete code repository $C=\{c_{i}\}_{i=1}^{n=|F|}$. We note that this iterative formulation can ensure that $i$-th code is generated with full awareness of its dependencies and the evolving state of the repository.
4 Experimental Setup
We now describe the experimental setup, including datasets, methods, and evaluation metrics.
4.1 Datasets
To validate the efficacy of our PaperCoder framework in transforming machine learning papers into their corresponding repository-level codes, we construct a new benchmark dataset, which we refer to as Paper2CodeBench. Specifically, we collect the accepted papers from recent machine learning venues (such as ICLR, ICML, and NeurIPS 2024) with the OpenReview API [^1] https://docs.openreview.net/reference/api-v2, and filter them based on the availability of code with its total number of tokens less than 70,000, to ensure the full repository remains within reasonable processing limits of modern LLMs for generation and evaluation. Also, to maintain the benchmark quality, we perform model-based evaluation [22] with GPT-4o on all the collected repositories and select the top 30 from each venue, resulting in a total of 90 papers listed in Tables 1112, and 13. Also, we additionally consider 21 papers for human evaluation (See Table 15). In addition to Paper2CodeBench, we also use the recently released PaperBench Code-Dev [38], which consists of 20 papers from ICML 2024 with paper-specific rubrics annotated by humans. In particular, those rubrics are used to judge the correct implementation based on LLM-based evaluation.
4.2 Baselines and Our Model
We target the novel problem of Paper2Code, and there are no baselines designed for it to enable direct comparison. Nevertheless, we consider several related approaches proposed to implement repository-level code (or the entire software) from natural language inputs (such as software requirements), in addition to the ablated variants of our full PaperCoder framework, as follows: ChatDev [34] is a multi-agent framework for software development, where several role-specific LLM-powered agents collaborate via structured dialogues; MetaGPT [13] similarity adopts a role-based multi-agent paradigm, but its development process is organized by the principle of Standardized Operating Procedures (SOPs); Abstract is a naive variant of our PaperCoder, which uses only the paper abstract for implementation; Paper, while using the full paper, prompts LLMs only once for code generation; PaperCoder (Ours) is our full framework, structured into three stages of planning, analysis, and code generation. In addition to them, for the PaperBench Code-Dev, we consider baselines suggested by it: Basic Agent is the agentic architecture that can run a predefined set of tools with the ReAct-style approach [49], built upon the basic agent from Inspect AI [^2] https://inspect.ai-safety-institute.org.uk/agents.html#sec-basic-agent, and Iterative Agent that extends Basic Agent, iteratively instructing the model to complete the next subtask.
4.3 Evaluation Setup
Recall that, as shown in Figure 1, the official code implementations of many papers are not available; however, manually annotating their corresponding code implementations to evaluate the quality of automatically generated code repositories is highly labor-intensive and challenging. To address this and ultimately perform the evaluation at scale, we design two model-based evaluation protocols: reference-based (when ground-truth code is available) and reference-free (when it is not), following the recent trends in using LLMs to judge the quality of model outputs [52, 11, 22], especially when they are long. In addition to this, we also perform human evaluations with the authors of the original papers, to ensure reliable judgments and to assess the quality of our model-based evaluations by measuring their correlation with human scores. We discuss each evaluation protocol in detail below.
Reference-Based Evaluation
In this evaluation setup, we use the official author-released repository as the gold standard if it is available, since it most accurately reflects the implementations intended by the authors, including the components they consider essential for reproducing the main results presented in the paper. Specifically, we prompt the evaluation model (such as o3-mini-high [^3] Unless otherwise stated, we use o3-mini-high due to strong code understanding and reasoning capability.) to judge the quality of the generated repository with respect to the gold repository, alongside the input paper as context (See Appendix E for the detailed prompt). During evaluation, the model then identifies key components (to be implemented), categorizes them into three severity levels (high, medium, and low), and critiques how well each component is implemented. After that, it returns the overall score on a 5-point Likert scale. We note that, to ensure the reliability of the model-based evaluation, we sample the outputs multiple times (for example, 8 generations per evaluation) and report the averaged score.
Reference-Free Evaluation
For cases where the official author-released code is not available, we introduce the reference-free evaluation protocol that leverages only the paper to assess the quality of its generated repository. In other words, similar to the reference-based evaluation protocol, the evaluation model is prompted to identify key components, categorize them by severity, and critique their implementations in the generated code, but they are performed solely based on the information provided in the paper without access to the ground-truth repository. The rest of the evaluation process, such as sampling and score averaging, follows the same setup as in the reference-based evaluation.
Human Evaluation
While model-based evaluation offers a scalable and automated way of assessment, we also conduct human evaluations to validate our PaperCoder based on the expert-grounded evaluation. Specifically, to ensure informed and accurate judgment, each participant is assigned a paper for which they are the first author. Furthermore, they are presented with multiple implementations generated by different approaches, and asked to rank them. We provide more details in Appendix A.2.
Evaluation on PaperBench Code-Dev
For evaluation on PaperBench Code-Dev [38], we follow their evaluation setup, measuring the score over the paper-specific rubrics with LLM-based evaluation.
Table 1: Results on our Paper2CodeBench, where we report average scores and standard deviations (in parentheses) grouped by conferences. Oracle denotes the evaluation results with the official repository released by the paper authors. Also, on the right side, we report statistics on the number of tokens, files, and functions, averaged over all implementations. Bold indicates the best scores, statistically significant than baselines ($p\leq 0.05$).
Reference-Based Evaluation Reference-Free Evaluation Statistics
ICLR ICML NeurIPS ICLR ICML NeurIPS # of Tokens # of Files # of Funcs
ChatDEV 2.70 (0.63) 2.97 (0.58) 2.96 (0.69) 4.00 (0.65) 4.12 (0.53) 4.01 (0.74) 6150.54 6.99 23.82
MetaGPT 2.48 (0.48) 2.75 (0.70) 2.95 (0.87) 3.52 (0.60) 3.63 (0.75) 3.59 (0.92) 5405.21 3.24 18.08
Abstract 2.28 (0.42) 2.43 (0.49) 2.35 (0.62) 3.03 (0.64) 3.01 (0.60) 2.99 (0.78) 3376.99 1.28 12.62
Paper 3.08 (0.66) 3.28 (0.67) 3.22 (0.80) 4.15 (0.63) 4.30 (0.53) 4.08 (0.84) 3846.33 1.79 14.84
PaperCoder 3.68 (0.52) 3.72 (0.54) 3.83 (0.50) 4.73 (0.32) 4.73 (0.44) 4.77 (0.38) 14343.38 6.97 35.22
Oracle N/A N/A N/A 4.84 (0.26) 4.80 (0.32) 4.83 (0.38) 32149.04 28.00 122.03
5 Experimental Results and Analysis
Main Results
Table 1 presents main results on Paper2CodeBench, from which we observe that PaperCoder consistently outperforms all baselines on both the reference-based and reference-free evaluation settings by substantial margins. We hypothesize that this performance gap of the proposed PaperCoder stems from its top-down behavior, analyzing full papers thoughtfully before starting code generation, unlike prior approaches that typically follow the bottom-up strategy, which begins with and expands short requiremental descriptions (via role-playing or Standardized Operating Procedures). In other words, the top-down approach, operationalized within PaperCoder through the sequential process of planning, analysis, and coding, is effective in handling long-form scientific documents, which are often loosely structured from a software engineering perspective (See Table 6 for more analyses). Also, when compared to the non-comparable Oracle setting (which performs evaluations on the author-released code repositories), PaperCoder achieves performance that is on par, with only marginal differences without a statistically significant gap, demonstrating the effectiveness of our approach in faithfully implementing code whose quality is closer to the implementation by authors.
Correlation between Reference-Based and Reference-Free Evaluation
Recall that the reference-free evaluation protocol is designed for cases where the ground-truth repository is not available, and to investigate whether it works as a reliable proxy for the reference-based evaluation protocol, we measure their rank correlation on all samples from Paper2CodeBench. Then, as shown in Figure 3, there is a strong positive correlation between the reference-based and reference-free evaluations, achieving a Pearson correlation coefficient of $r=0.79$. This result supports that the reference-free evaluation (despite the absence of ground-truth repositories) can serve as a reliable proxy for the reference-based evaluation, ultimately functioning as a standalone metric to assess the code quality.
Figure 3: Correlation between model-based evaluations: reference-based and reference-free.
Table 2: Results with human evaluation. For model-based evaluations (both reference-based and reference-free), 5-point Likert evaluation scores are converted to rankings for comparability with human ranking results. Human rankings are also converted to scores of 5 (top repository), 3 (middle repository), and 1 (bottom repository).
Score ($\uparrow$) Ranking ($\downarrow$)
Ref-based Ref-free Human Ref-based Ref-free Human
Abstract 2.26 (0.37) 2.94 (0.61) 2.68 (0.56) 2.96 (0.20) 2.96 (0.00) 2.70 (0.56)
Paper 3.00 (0.54) 3.91 (0.63) 2.76 (1.20) 1.92 (0.41) 1.88 (0.38) 2.09 (0.60)
PaperCoder (Ours) 3.66 (0.43) 4.55 (0.51) 4.60 (1.00) 1.08 (0.28) 1.08 (0.28) 1.22 (0.52)
ChatDEV 2.68 (0.60) 3.82 (0.37) 2.12 (1.17) 2.58 (0.50) 2.23 (0.59) 2.43 (0.59)
MetaGPT 2.61 (0.54) 3.39 (0.67) 2.12 (1.17) 2.38 (0.58) 2.46 (0.51) 2.43 (0.59)
PaperCoder (Ours) 3.66 (0.43) 4.55 (0.51) 4.76 (0.88) 1.04 (0.20) 1.04 (0.20) 1.13 (0.46)
Table 3: PaperBench Code-Dev results. We report the averaged performance over three runs with standard deviations.
Replication Score (%)
Model o3-mini-high claude-3.5-sonnet
BasicAgent $5.1\pm 0.8$ $35.4\pm 0.8$
IterativeAgent $16.4\pm 1.4$ $27.5\pm 1.6$
PaperCoder $\textbf{45.14}\pm 0.3$ $\textbf{51.14}\pm 1.4$
Table 4: Results based on both model-based and human evaluations with varying backbone LLMs for PaperCoder.
DS-Coder Qwen-Coder DS-Distill-Qwen o3-mini-high
Score ($\uparrow$) Ref-based 1.47 (0.46) 1.78 (0.28) 2.05 (0.25) 3.66 (0.43)
Ref-free 1.62 (0.54) 2.09 (0.22) 2.31 (0.24) 4.55 (0.51)
Human 1.32 (0.58) 2.71 (1.12) 3.29 (0.98) 4.68 (0.80)
Ranking ($\downarrow$) Ref-based 3.46 (0.00) 2.92 (0.88) 2.25 (0.65) 1.00 (0.20)
Ref-free 3.50 (0.00) 2.88 (0.83) 2.12 (0.54) 1.00 (0.25)
Human 3.74 (0.45) 2.74 (0.86) 2.30 (0.70) 1.22 (0.60)
Table 5: Rank correlation coefficient between human and model-based evaluations (with GPT-4o or o3-mini).
GPT-4o o3-mini-high
Ref-based 0.74 0.78
Ref-free 0.71 0.73
Human Evaluation Results
In addition to automatic model-based evaluations, we conduct human evaluations and report the results in Table 2. From this, we confirm that PaperCoder achieves the best ranking among all compared methods, alongside the best scores with model-based evaluations, which reaffirms its effectiveness. Moreover, to ensure whether the model-based evaluations are a reasonable proxy to judge the implementation quality for papers, we measure their correlations with human evaluation scores. As shown in Table 5, we observe strong rank correlations across both reference-based and reference-free settings (where GPT-4o and o3-mini-high are used as evaluation models), which suggests that model-based evaluation can reliably approximate human judgment. Also, based on this result, we use o3-mini-high as the default evaluation model in our experiments due to its highest correlation. Lastly, we ensure the quality and reliability of human evaluations by measuring the inter-annotator agreement score based on Cohen’s kappa coefficient, which exhibits a high score of 0.79, indicating strong consistency.
Results on PaperBench Code-Dev
In addition to our Paper2CodeBench, we further validate the effectiveness of PaperCoder on another PaperBench Code-Dev dataset, which enables fine-grained evaluations for code implementations. As Table 3 shows, PaperCoder achieves the highest replication scores across two different LLMs of o3-mini-high and Claude 3.5 Sonnet, substantially outperforming baselines designed for PaperBench Code-Dev. These results further demonstrate the generalizability and robustness of PaperCoder across diverse evaluation benchmarks and models.
Analysis on Different LLMs
Extending the model variations results on PaperBench Code-Dev, we conduct an auxiliary analysis with DS-Coder (DeepSeek-Coder-V2-Lite-Instruct [8]), Qwen-Coder (Qwen2.5-Coder-7B-Instruct [15]), DS-Distill-Qwen (DeepSeek-R1-Distill-Qwen-14B [9]), and o3-mini-high (the high reasoning-effort variant of o3-mini) on Paper2CodeBench. As summarized in Table 4, the proprietary model (o3-mini-high) consistently outperforms all other backbones across all evaluation settings. Among other open-source models, DS-Distill-Qwen performs the best, followed by Qwen-Coder and DS-Coder. These results suggest the importance of selecting a capable backbone to instantiate PaperCoder, particularly one with strong reasoning capabilities. Also, based on this result, we primarily use o3-mini-high as the basis backbone model for all experiments.
Fine-Grained Analysis of Generated Repositories
To more thoroughly evaluate the quality and practical utility of the generated code, we conduct a set of fine-grained human analyses, according to two perspectives: its usability in reproducing the original work and its component-wise implementation quality. Specifically, we first ask the human annotators whether the top-ranked repository generated by PaperCoder would make it easier to reproduce the methods and experiments compared to starting from scratch. Then, a majority of participants (92%) agree that it would facilitate reproduction, reinforcing the real-world usefulness of our approach. Also, we conduct a component-level implementation analysis to assess which parts of the papers are most effectively translated into code, by asking human annotators to identify the key component-specific elements for Data Processing, Method, and Evaluation, and then measure how many of these elements are actually implemented in the generated code. As shown in Figure 5, we find that the sections of Method, Evaluation, and Data Processing achieve the highest coverage scores, in that order, with each exceeding 50%. Notably, among the errors observed, many of them stem from missing or incorrect implementation details, often rooted in uncertainties about the structure and format of the data, as well as how the data is loaded, which are often under-specified in the papers. Lastly, to investigate why human annotators prefer PaperCoder over its baselines and ablated variants (with 22 out of 25 selecting the repositories from PaperCoder), we ask them to provide the reasons for their choices, and the majority of which are completeness, clean structure, and faithfulness to the original papers, summarized in Table 10.
Ablation Studies
To see how much each component of PaperCoder contributes to the performance gain, we conduct ablation studies on the subset of Paper2CodeBench (composed of ICML papers). Specifically, we start with the base method that uses only the full paper (called Paper) and incrementally add components in the order they are executed (such as overall plan, architecture design, logic design, configuration generation, and final analysis), whose results are reported in Table 6. From this, we observe that the performance steadily improves as additional components are incorporated. Also, we observe a performance drop when the architecture design module is added; however, while this might seem surprising at first, it is in fact expected: architecture design alone does not specify the execution or implementation order of files, which leads to confusion during the code generation stage. However, this issue is effectively addressed once the subsequent logic design module explicitly defines file dependencies and establishes a clear generation order, leading to a substantial performance boost. Overall, integrating all modules in the pipeline yields the highest performance, confirming the effectiveness of our fully structured, multi-stage pipeline, with each module playing a valuable role.
Correlation on Paper Type
To see whether the acceptance category (or presentation format) of papers correlates with the quality of their corresponding implementations by PaperCoder, we analyze it by separating papers into oral/spotlight and poster categories on Paper2CodeBench (which includes 14 oral or spotlight papers and 76 poster papers). As shown in Figure 5, scores are slightly higher for oral/spotlight papers on model-based evaluations with GPT-4o and o3-mini, suggesting that papers with higher recognition might reflect clearer writing, probably leading to faithful code generation.
Figure 4: Detailed, fine-grained analyses on code by PaperCoder.
Table 6: Ablation results on the subset of Paper2CodeBench (ICML papers) with scores and standard deviations.
Ref-based Ref-free
Abstract 2.43 (0.49) 3.01 (0.60)
Paper 3.28 (0.67) 4.30 (0.53)
+ Overall Plan 3.40 (0.57) 4.34 (0.58)
+ Arch. Design 3.13 (0.68) 4.07 (0.74)
+ Logic Design 3.60 (0.52) 4.50 (0.57)
+ Config File 3.66 (0.45) 4.45 (0.53)
+ Analysis (Ours) 3.72 (0.54) 4.73 (0.44)
Figure 5: Model-based evaluation results by presentation type of papers (poster vs. oral/spotlight).
Analysis on Executability
It is worth noting that making the repository-level code executable and fully reproducible in one go is extremely challenging (even for human developers), as demonstrated by prior findings from Starace et al. [38]. Also, our goal is to provide a strong and faithful starting point that meaningfully aids reproduction efforts (supported by our results), rather than aiming for perfect reproduction out of the box. Nevertheless, to assess how close our generated repositories are to being directly executable, we perform manual execution evaluations on five papers. Specifically, when execution errors occur, we manually debug and refine the code and adapt the input data as needed to enable successful runs. We then find that, on average, only 0.81% of the code lines require minor modification, such as updating deprecated API calls or correcting data type mismatches, for successful execution (See Examples in Figures 6 and 7, alongside the detailed statistics in Table 9), which highlights that our generated repositories are near-executable with minimal human intervention.
6 Conclusion
In this work, we introduced PaperCoder, a framework that automatically generates code repositories from research papers in machine learning through a structured, three-stage pipeline. Specifically, we defined a high-level roadmap, system architecture, execution logic, and configuration via the planning stage, which are then enhanced through detailed per-file analysis, followed by the sequential code generation informed by artifacts from prior stages. To validate PaperCoder, we performed evaluations on two benchmarks: our Paper2CodeBench, comprising recent papers from top-tier machine learning venues, and (recently released) PaperBench Code-Dev, providing fine-grained evaluation protocols, on which PaperCoder consistently outperforms existing baselines on both model-based and human evaluations. Furthermore, additional analyses demonstrate its robustness and practicality: it remains effective across different LLM backbones, shows strong executability with only 0.81% of the lines requiring minor fixes, and benefits from each stage in the pipeline. We envision PaperCoder as one important step toward accelerating scientific progress by aiding the reproduction of research papers.
Appendix A Additional Experimental Designs
A.1 Implementation Details
All experiments are conducted using the o3-mini with high reasoning effort version (o3-mini-high) as the default backbone, released on January 31, 2025. To collect paper metadata and content, we use openreview_scraper [^4] https://github.com/pranftw/openreview_scraper with the OpenReview API [^5] https://docs.openreview.net/reference/api-v2 and the Semantic Scholar API [^6] https://www.semanticscholar.org/product/api. For document processing, we convert papers into structured JSON format using the s2orc-doc2json library [23] [^7] https://github.com/allenai/s2orc-doc2json. Notably, with o3-mini-high to generate repositories for 90 papers, the total API cost of PaperCoder amounts to \$76.65, resulting in an average cost of approximately \$0.90 per paper.
A.2 Human Evaluation Process
Given the complexity of the task (requiring comprehension of scientific papers and their associated implementations), we recruit participants who have at least one peer-reviewed paper and a degree in computer science. We note that they were compensated at a rate of \$15 per hour. For annotation, they were provided with a 4-page document, which includes task instructions, annotation examples, and 10 generated repositories grouped into three sets, as follows: (Group 1) Model Variants of Our Method that includes repositories generated by our system using different backbone models (e.g., o3-mini vs. three open-source alternatives); (Group 2) Naive Baselines that includes repositories generated using only the Paper or the Abstract as input; and (Group 3) Related Works that includes repositories generated by existing software development frameworks, such as MetaGPT and ChatDev. Each repository was anonymized using a repo X naming format to prevent bias regarding the generation method. Following the question guidelines in the document, annotators reviewed and evaluated the repositories generated by different methods and models. Also, on average, evaluating 10 repositories for a single paper took approximately 45 minutes. Table 23 shows a detailed annotation example.
A.3 Reference-Based Evaluation
In the reference-based evaluation setup, the repository may exceed the context length of (even frontier) LLMs. Following Starace et al. [38], when this occurs, we prompt the model to select the most relevant files for evaluation. The selected subset is then used as the reference for scoring. We use the gpt-4o-2024-11-20 as the evaluation model.
A.4 PaperBench Code-Dev Evaluation
While PaperCoder is designed to generate only the source code, the PaperBench Code-Dev benchmark used for evaluation requires an additional script file called reproduce.sh. To meet this requirement, we further prompt the coding agent to generate it for this benchmark and evaluate the code with it.
Appendix B Additional Experimental Results and Analysis
B.1 Code Availability
To estimate the proportion of accepted papers that release official code repositories, we collect data from three major machine learning conferences in 2024: ICLR, ICML, and NeurIPS. Specifically, we first retrieve the list of accepted papers from each conference using the OpenReview API [^8] https://docs.openreview.net/reference/api-v2 via openreview_scraper [^9] https://github.com/pranftw/openreview_scraper. While OpenReview abstracts sometimes include repository links, they are more commonly found in ArXiv [^10] https://arxiv.org/ abstracts. Therefore, we additionally use the Semantic Scholar API [^11] https://www.semanticscholar.org/product/api to obtain ArXiv abstracts corresponding to the accepted papers. We then check whether the abstract includes a GitHub URL as an indicator of released code. Table 7 summarizes the number of accepted papers, the number with publicly available repositories, and the corresponding percentages for each conference. On average, only 19.5% of accepted papers in them provide official code.
Table 7: Code availability across major machine learning conferences. We report the total number of accepted papers, the number of papers with publicly available code (identified via GitHub URLs in ArXiv abstracts), and the corresponding percentage for each venue. The last row shows the average across all three conferences.
Conference # of Accepted w/ Code Percentage (%)
ICLR 2024 2207 467 21.2
ICML 2024 2610 435 16.7
NeurIPS 2024 4006 825 20.6
Average 2941 576 19.5
Table 8: Average Replication Scores (%) on PaperBench Code-Dev. For all OpenAI models, the reasoning effort is set to high, and we take results for BasicAgent and IterativeAgent from the PaperBench paper [38]. For PaperCoder, we report the average and standard deviation over three runs, except for o1 and o3 due to costs.
Model Replication Score (%) Cost per Paper (\$)
BasicAgent (o3-mini) $5.1\pm 0.8$ N/A
BasicAgent (o1) $19.5\pm 1.2$ N/A
BasicAgent (claude-3-5-sonnet) $35.4\pm 0.8$ N/A
IterativeAgent (o3-mini) $16.4\pm 1.4$ N/A
IterativeAgent (o1) $43.3\pm 1.1$ 400.00
IterativeAgent (claude-3-5-sonnet) $27.5\pm 1.6$ N/A
PaperCoder (o3-mini) $45.14\pm 0.3$ 0.69
PaperCoder (o1) $38.31$ 8.81
PaperCoder (o3) $60.86$ 8.99
PaperCoder (claude-3-5-sonnet) $51.14\pm 1.4$ 3.61
B.2 PaperBench Code-Dev Results
We conduct additional experiments using various reasoning models, as shown in Table 8. Overall, our method achieves strong replication scores across models. Notably, when using o3, PaperCoder records the highest score of 60.86%. These results suggest that the latest and larger models, particularly those with stronger reasoning and coding capabilities, tend to yield better performance.
Appendix C Limitations and Future Work
While PaperCoder demonstrates strong performance in reproducing machine learning papers (where code implementations are particularly helpful and usually necessary for validating research ideas), its current scope is limited to this domain. Beyond this, we believe accelerating the reproduction of scientific discovery to other domains where code is not the primary medium for validation, such as theoretical mathematics, is an exciting direction for future work. In addition, the current version of PaperCoder processes only textual inputs, and extending it to process visual inputs (such as figures in papers) is an interesting avenue. Lastly, as with other repository-level code generation approaches, improving executability remains an important (but still challenging) direction for future work.
Appendix D Broader Impacts and Ethics Statement
Our work aims to generate faithful code repositories from scientific papers in machine learning, and we believe it has a substantial positive impact in contributing to open science and facilitating rapid experimentation. However, we also acknowledge potential risks and misuse of our framework. For example, some papers intentionally refrain from releasing implementations due to security concerns, such as those involving jailbreaking or exploitation techniques. Yet, our method could potentially be used to reproduce such sensitive implementations. To address such risks, in real-world production, it would be necessary to develop and incorporate safeguards (such as harmful content filters, protective prompting, and secure execution environments) to ensure responsible and safe use of our framework.
Table 9: Executability results across generated repositories. We manually select five papers and generate corresponding repositories using PaperCoder. For each repository, we report the number of lines modified during debugging, the total number of lines, and the percentage of modified lines.
Repo Name CoLoR cognitive-behaviors RADA Self-Instruct G-EVAL Average
Modified lines (*.py) 2 0 10 26 10 8
Modified lines (config.yaml) 3 6 7 1 4 3.5
Total lines 1132 2060 1609 1334 1374 1251.5
Percentage 0.44 0.29 1.06 2.02 1.02 0.81
Table 10: Qualitative analysis of top-ranked repositories. We categorize the reasons why human annotators select the repositories generated by our framework as their top choice into six (described in the first row).
Completeness Clean Structure Faithfulness to Paper Ease of Use Code Quality Unique Strengths
16 13 8 6 7 4
Original 1# data_loader.py 2import openai 3import pandas as pd 4from rouge_score import rouge_scorer 5 6# Set up basic logging configuration 7logging.basicConfig(level=logging.INFO, format="%(asctime)s [%(levelname)s] %(message)s") 8 9 10...(omited)... 11 12 response = openai.Completion.create( 13 engine=self.engine, 14 prompt=prompt, 15 max_tokens=150, 16 temperature=0.7, 17 n=1, 18 stop=["Task 16"] 19 ) 20 raw_text = response.choices[0].text.strip()
Editied 1# data_loader.py 2from openai import OpenAI 3import pandas as pd 4from rouge_score import rouge_scorer 5 6# Set up basic logging configuration 7logging.basicConfig(level=logging.INFO, format="%(asctime)s [%(levelname)s] %(message)s") 8 9openai_api_key = os.environ.get("OPENAI_API_KEY") 10client = OpenAI(api_key=openai_api_key) 11 12...(omited)... 13 14 response = client.chat.completions.create( 15 model=self.engine, 16 messages=[{"role": "user", "content": prompt}], 17 max_completion_tokens=150, 18 temperature=0.7, 19 n=1, 20 stop=["Task 16"] 21 ) 22 raw_text = response.choices[0].message.content.strip()
Figure 6: Error case from Self-Instruct. The left shows the original code generated by PaperCoder using o3-mini-high, and the right shows the version manually edited by the authors to correct the error. In this example, an outdated API call is updated to its latest version. In the original version, lines 2, 12, 14, 15, and 29 were removed; in the edited version, lines 2, 9, 10, 14, 16, 17, and 22 were added.
Original 1# trainer.py 2self.optimizer = AdamW(self.model.model.parameters(), lr=lr) 3 4# model.py 5self.model = AutoModelForCausalLM.from_pretrained(base_model)
Edited 1# trainer.py 2self.optimizer = AdamW(self.model.model.parameters(), lr=float(lr)) 3 4# model.py 5self.model = AutoModelForCausalLM.from_pretrained(base_model, trust_remote_code=True)
Figure 7: Error case from CoLoR. The left shows the original code generated by PaperCoder using o3-mini-high, and the right shows the manually edited version by the authors. In this example, a numeric value is cast correctly and a required argument is added to enable execution. Lines 2 and 5 are modified in the corrected version.
Appendix E Prompts
Prompt for generating the overall plan in the planning stage [System] You are an expert researcher and strategic planner with a deep understanding of experimental design and reproducibility in scientific research. You will receive a research paper in JSON format. Your task is to create a detailed and efficient plan to reproduce the experiments and methodologies described in the paper. This plan should align precisely with the paper’s methodology, experimental setup, and evaluation metrics. Instructions: 1. Align with the Paper: Your plan must strictly follow the methods, datasets, model configurations, hyperparameters, and experimental setups described in the paper. 2. Be Clear and Structured: Present the plan in a well-organized and easy-to-follow format, breaking it down into actionable steps. 3. Prioritize Efficiency: Optimize the plan for clarity and practical implementation while ensuring fidelity to the original experiments. [User] ## Paper {paper_json} ## Task 1. We want to reproduce the method described in the attached paper. 2. The authors did not release any official code, so we have to plan our own implementation. 3. Before writing any Python code, please outline a comprehensive plan that covers: - Key details from the paper’s **Methodology**. - Important aspects of **Experiments**, including dataset requirements, experimental settings, hyperparameters, or evaluation metrics. 4. The plan should be as **detailed and informative** as possible to help us write the final code later. ## Requirements - You don’t need to provide the actual code yet; focus on a **thorough, clear strategy**. - If something is unclear from the paper, mention it explicitly. ## Instruction The response should give us a strong roadmap, making it easier to write the code later.
Figure 8: Prompt for generating the overall plan in the planning stage.
[…]
Figure 9: Prompt for generating the architecture design in the planning stage. This prompt follows the previous prompt and response shown in Figure 8.
Prompt for generating the logic design in the planning stage [User] Your goal is break down tasks according to PRD/technical design, generate a task list, and analyze task dependencies. You will break down tasks, analyze dependencies. You outline a clear PRD/technical design for reproducing the paper’s method and experiments. Now, let’s break down tasks according to PRD/technical design, generate a task list, and analyze task dependencies. The Logic Analysis should not only consider the dependencies between files but also provide detailed descriptions to assist in writing the code needed to reproduce the paper. —– ## Format Example [CONTENT] { "Required packages": [ "numpy==1.21.0", "torch==1.9.0" ], "Required Other language third-party packages": [ "No third-party dependencies required" ], "Logic Analysis": [ [ "data_preprocessing.py", "DataPreprocessing class …….." ], [ "trainer.py", "Trainer ……. " ], [ "dataset_loader.py", "Handles loading and …….." ], [ "model.py", "Defines the model ……." ], [ "evaluation.py", "Evaluation class …….. " ], [ "main.py", "Entry point ……." ] ], "Task list": [ "dataset_loader.py", "model.py", "trainer.py", "evaluation.py", "main.py" ], "Full API spec": "openapi: 3.0.0 …", "Shared Knowledge": "Both data_preprocessing.py and trainer.py share ……..", "Anything UNCLEAR": "Clarification needed on recommended hardware configuration for large-scale experiments." } [/CONTENT] ## Nodes: "<node>: <type> # <instruction>" - Required packages: typing.Optional[typing.List[str]] # Provide required third-party packages in requirements.txt format.(e.g., ’numpy==1.21.0’). - Required Other language third-party packages: typing.List[str] # List down packages required for non-Python languages. If none, specify "No third-party dependencies required". - Logic Analysis: typing.List[typing.List[str]] # Provide a list of files with the classes/methods/functions to be implemented, including dependency analysis and imports. Include as much detailed description as possible. - Task list: typing.List[str] # Break down the tasks into a list of filenames, prioritized based on dependency order. The task list must include the previously generated file list. - Full API spec: <class ’str’> # Describe all APIs using OpenAPI 3.0 spec that may be used by both frontend and backend. If front-end and back-end communication is not required, leave it blank. - Shared Knowledge: <class ’str’> # Detail any shared knowledge, like common utility functions or configuration variables. - Anything UNCLEAR: <class ’str’> # Mention any unresolved questions or clarifications needed from the paper or project scope. ## Constraint Format: output wrapped inside [CONTENT][/CONTENT] like the format example, nothing else. ## Action Follow the node instructions above, generate your output accordingly, and ensure it follows the given format example.
Figure 10: Prompt for generating the logic design in the planning stage. This prompt follows the previous prompt and response shown in Figure 9.
Prompt for generating the configuration file in the planning stage [User] You write elegant, modular, and maintainable code. Adhere to Google-style guidelines. Based on the paper, plan, design specified previously, follow the "Format Example" and generate the code. Extract the training details from the above paper (e.g., learning rate, batch size, epochs, etc.), follow the "Format example" and generate the code. DO NOT FABRICATE DETAILS — only use what the paper provides. You must write ‘config.yaml‘. ATTENTION: Use ’##’ to SPLIT SECTIONS, not ’#’. Your output format must follow the example below exactly. —– # Format Example ## Code: config.yaml “‘yaml ## config.yaml training: learning_rate: … batch_size: … epochs: … “‘ —– ## Code: config.yaml
Figure 11: Prompt for generating the configuration file in the planning stage. This prompt follows the previous prompt and response shown in Figure 10.
Prompt for analysis [System] You are an expert researcher, strategic analyzer and software engineer with a deep understanding of experimental design and reproducibility in scientific research. You will receive a research paper in JSON format, an overview of the plan, a design in JSON format consisting of "Implementation approach", "File list", "Data structures and interfaces", and "Program call flow", followed by a task in JSON format that includes "Required packages", "Required other language third-party packages", "Logic Analysis", and "Task list", along with a configuration file named "config.yaml". Your task is to conduct a comprehensive logic analysis to accurately reproduce the experiments and methodologies described in the research paper. This analysis must align precisely with the paper’s methodology, experimental setup, and evaluation criteria. 1. Align with the Paper: Your analysis must strictly follow the methods, datasets, model configurations, hyperparameters, and experimental setups described in the paper. 2. Be Clear and Structured: Present your analysis in a logical, well-organized, and actionable format that is easy to follow and implement. 3. Prioritize Efficiency: Optimize the analysis for clarity and practical implementation while ensuring fidelity to the original experiments. 4. Follow design: YOU MUST FOLLOW "Data structures and interfaces". DONT CHANGE ANY DESIGN. Do not use public member functions that do not exist in your design. 5. REFER TO CONFIGURATION: Always reference settings from the config.yaml file. Do not invent or assume any values—only use configurations explicitly provided. [User]
# Context
## Paper {The content of the paper in json format} —– ## Overview of the plan {The content of the overall plan} —– ## Design {The content of the architecture design} —– ## Task {The content of the logic design} —– ## Configuration file “‘yaml {The content of the configuration file} “‘ —– ## Instruction Conduct a Logic Analysis to assist in writing the code, based on the paper, the plan, the design, the task and the previously specified configuration file (config.yaml). You DON’T need to provide the actual code yet; focus on a thorough, clear analysis. Write the logic analysis in ’{The name of the file to be generated"}’, which is intended for ’{Description of the file generated through the "Logic Analysis" step of the logic design.}’. —– ## Logic Analysis: {todo_file_name}
Figure 12: Prompts for analysis. {} indicate placeholders to be filled with the content described in the accompanying explanation. The prompt is presented to the LLM for each file, following the sequence defined in the logic design.
Prompt for coding [System] You are an expert researcher and software engineer with a deep understanding of experimental design and reproducibility in scientific research. You will receive a research paper in JSON format, an overview of the plan, a Design in JSON format consisting of "Implementation approach", "File list", "Data structures and interfaces", and "Program call flow", followed by a Task in JSON format that includes "Required packages", "Required other language third-party packages", "Logic Analysis", and "Task list", along with a configuration file named "config.yaml". Your task is to write code to reproduce the experiments and methodologies described in the paper. The code you write must be elegant, modular, and maintainable, adhering to Google-style guidelines. The code must strictly align with the paper’s methodology, experimental setup, and evaluation metrics. Write code with triple quoto. [User]
# Context
## Paper {The content of the paper in json format} —– ## Overview of the plan {The content of the overall plan} —– ## Design {The content of the architecture design} —– ## Task {The content of the logic design} —– ## Configuration file “‘yaml {The content of the configuration file} “‘ —– ## Code Files {The content of the code files generated in the previous step.} —– # Format example ## Code: {todo_file_name} “‘python ## todo_file_name “‘ —– # Instruction Based on the paper, plan, design, task and configuration file(config.yaml) specified previously, follow "Format example", write the code. We have {done_file_lst}. Next, you must write only the "{todo_file_name}". 1. Only One file: do your best to implement THIS ONLY ONE FILE. 2. COMPLETE CODE: Your code will be part of the entire project, so please implement complete, reliable, reusable code snippets. 3. Set default value: If there is any setting, ALWAYS SET A DEFAULT VALUE, ALWAYS USE STRONG TYPE AND EXPLICIT VARIABLE. AVOID circular import. 4. Follow design: YOU MUST FOLLOW "Data structures and interfaces". DONT CHANGE ANY DESIGN. Do not use public member functions that do not exist in your design. 5. CAREFULLY CHECK THAT YOU DONT MISS ANY NECESSARY CLASS/FUNCTION IN THIS FILE. 6. Before using a external variable/module, make sure you import it first. 7. Write out EVERY CODE DETAIL, DON’T LEAVE TODO. 8. REFER TO CONFIGURATION: you must use configuration from "config.yaml". DO NOT FABRICATE any configuration values. {detailed_logic_analysis} ## Code: {todo_file_name}
Figure 13: Prompts for coding. {} indicate placeholders to be filled with the content described in the accompanying explanation. The prompt is presented to the LLM for each file, following the sequence defined in the logic design. Previously generated code files are accumulated and provided as part of the ## Code Files input.
Prompt for model-based reference-based evaluation [System] You will be given a research paper along with two corresponding code repositories: a gold repository and a target repository. Your task is to compare the target repository against the gold repository, rate the target repository on one metric, and provide a critique highlighting key differences. Please make sure you read and understand these instructions carefully. Keep this document open while reviewing, and refer to it as needed. Evaluation Criteria: Correctness (1-5): The quality of the target repository in accurately implementing the paper’s concepts, methodology, and algorithms without logical errors, as compared to the gold repository. Additionally, provide a critique focusing on the completeness, accuracy, and implementation choices made in the target repository relative to the gold repository. 1: Very Poor. The target repository does not correctly implement the core concepts, methodology, or algorithms from the paper. Major logical errors or missing components are present, especially when compared to the gold repository. 2: Poor. The target repository attempts to implement the paper’s concepts but contains significant mistakes or missing components, making the implementation incorrect when compared to the gold repository. 3: Fair. Some core components and concepts are correctly implemented in the target repository, but there are notable logical errors or inaccuracies compared to the gold repository. 4: Good. The target repository correctly implements the key components and methodology, with only minor inaccuracies or deviations from the gold repository. 5: Excellent. The target repository fully and accurately implements all relevant key components, methodology, and algorithms from the paper, matching the quality of the gold repository. Evaluation Steps 1. Identify Key Aspects of the Paper: Carefully read the research paper to understand its core concepts, methodology, and algorithms. Pay close attention to the key aspects that are crucial for implementing the paper’s results (e.g., specific algorithms, data preprocessing steps, evaluation protocols). 2. Analyze the Gold Repository: Examine the gold repository to understand how these key aspects have been implemented. Use the gold repository as a reference for how the paper’s methodology should be translated into code. Note the completeness, accuracy, and design choices in the gold repository that faithfully represent the paper’s concepts. 3. Examine the Target Repository: Analyze the target repository to assess how well it implements the key aspects of the paper. Reference the gold repository as a guide for understanding these key aspects in the target repository. Focus on whether the target repository’s core logic, algorithms, and structure align with the methodology and experiments described in the paper. 4. Identify Logical Errors and Deviations: Check for logical errors, missing steps, or deviations from the paper’s methodology. Note any incorrect representations, inconsistencies, or incomplete implementations that could affect the correctness of the target repository. 5. Provide a Critique: Consider both the completeness and accuracy of the implementation relative to the paper’s goals and the gold repository’s standard. You do not need to analyze minor details like logging functions, script organization, or documentation quality. Instead, concentrate on the correctness of the logic and implementation that ensures the core concepts from the paper are fully reflected in the target repository. For each mismatch or deviation in implementation, note down specific critiques comparing relevant functions in the target repository to the corresponding functions in the gold repository. Highlight incorrect logic, missing steps, or deviations that affect the correct implementation of the paper’s methodology. 5. Assess the Correctness: Determine whether the target repository includes all the critical elements described in the paper and implemented in the gold repository. Identify missing components, significant deviations, or incorrect implementations that could affect the correctness of the target repository. 6. Assign a Score: Based on your evaluation, provide a critique and assign a correctness score from 1 to 5 for the target repository, reflecting how well it implements the key aspects of the paper refer to the gold repository. Include a detailed critique in the specified JSON format. Severity Level: Each identified critique will be assigned a severity level based on its impact on the correctness of the methodology implementation. - High: Missing or incorrect implementation of the paper’s core concepts, major loss functions, or experiment components that are fundamental to reproducing the paper’s methodology. - Example: The main algorithm is missing or fundamentally incorrect. - Medium: Issues affecting training logic, data preprocessing, or other core functionalities that significantly impact performance but do not completely break the system. - Example: Improper training loop structure, incorrect data augmentation, or missing essential components in data processing. - Low: Errors in specific features that cause deviations from expected results but can be worked around with modifications. Any errors in the evaluation process belong to this category unless they impact the core concepts. These include minor issues like logging, error handling mechanisms, configuration settings, evaluation steps that do not alter the fundamental implementation and additional implementations not explicitly stated in the paper. - Example: Suboptimal hyperparameter initialization, incorrect learning rate schedule, inaccuracies in evaluation metrics, using a different random seed, variations in batch processing, different weight initialization, issues in result logging or reporting, variations in evaluation dataset splits, improper error handling in non-critical steps, mismatches in secondary evaluation criteria, or additional implementation details not specified in the paper that do not interfere with core results.
Prompt for model-based reference-based evaluation Example JSON format: “‘json { "critique_list": [ { "gold_file_name": "preprocessing.py", "gold_func_name": "data_process", "target_file_name": "dataset.py", "target_func_name": "train_preprocess", "severity_level": "medium", "critique": "A critique of the target repository’s file with reference to the gold repository." }, { "gold_file_name": "utils.py", "gold_func_name": "calculate_metric", "target_file_name": "metric.py", "target_func_name": "f1_at_k" "severity_level": "low", "critique": "A critique of the target repository’s file with reference to the gold repository." }, ], "score": 2 } “‘ Sample: Research Paper: {{The content of the paper}} Gold Repository: {{The gold repository, officially released by the authors, serves as the reference implementation.}} Target Repository: {{The generated repository, which serves as the target repository for evaluation.}} Please provide critique of the target repository and a single numerical rating (1, 2, 3, 4, or 5) based on the quality of the sample, following the Example JSON format, without any additional commentary, formatting, or chattiness.
Figure 14: Prompts for model-based reference-based evaluation. {{}} indicate placeholders to be filled with the content described in the accompanying explanation.
Prompt for model-based reference-free evaluation [System] You will be given a research paper along with its corresponding code repository. Your task is to rate the code repository on one metric and provide a critique highlighting key differences. Please make sure you read and understand these instructions carefully. Keep this document open while reviewing, and refer to it as needed. Evaluation Criteria: Correctness (1-5): The quality of the repository in accurately implementing the paper’s concepts, methodology, and algorithms without logical errors. Additionally, provide a critique focusing on the completeness, accuracy, and implementation choices made in the repository relative to the methodology and algorithms described in the paper. 1: Very Poor. The repository does not correctly implement the core concepts, methodology, or algorithms from the paper. Major logical errors or missing components are present. 2: Poor. The repository attempts to implement the paper’s concepts but contains significant mistakes or missing components, making the implementation incorrect. 3: Fair. Some core components and concepts are correctly implemented, but there are notable logical errors or inaccuracies in the methodology. 4: Good. The repository correctly implements the key components and methodology, with only minor inaccuracies that do not significantly affect correctness. 5: Excellent. The repository fully and accurately implements all key components, methodology, and algorithms from the paper without logical errors. Evaluation Steps 1. Identify Key Aspects of the Paper: Carefully read the paper to understand its core concepts, methodology, and algorithms. Pay close attention to key aspects crucial for implementing the paper’s results (e.g., specific algorithms, data preprocessing steps, evaluation protocols). 2. Examine the Code Repository: Analyze the repository to determine how well it implements the key aspects of the paper. Focus on whether the repository’s core logic, algorithms, and structure align with the methodology and experiments described in the paper. 3. Identify Logical Errors and Deviations: Check for logical errors, missing steps, or deviations from the paper’s methodology. Note any incorrect representations, inconsistencies, or incomplete implementations that could affect the correctness of the repository. 4. Provide a Critique: Consider the completeness and accuracy of the implementation relative to the paper’s goals. You do not need to analyze minor details like logging functions, script organization, or documentation quality. Instead, concentrate on the correctness of the logic and implementation to ensure the core concepts from the paper are fully reflected in the repository. For each identified issue, write a detailed critique specifying the affected files and functions in the repository. Highlight missing or incorrectly implemented steps that impact the correctness and alignment with the paper’s methodology. 5. Assess Completeness and Accuracy: Evaluate the repository for its completeness and accuracy relative to the paper’s methodology. Ensure that all critical components—such as data preprocessing, core algorithms, and evaluation steps—are implemented and consistent with the paper’s descriptions. 6. Assign a Score: Based on your evaluation, provide a critique and assign a correctness score from 1 to 5 for the repository, reflecting how well it implements the key aspects of the paper. Include a detailed critique in the specified JSON format. Severity Level: Each identified critique will be assigned a severity level based on its impact on the correctness of the methodology implementation. - High: Missing or incorrect implementation of the paper’s core concepts, major loss functions, or experiment components that are fundamental to reproducing the paper’s methodology. - Example: The main algorithm is missing or fundamentally incorrect. - Medium: Issues affecting training logic, data preprocessing, or other core functionalities that significantly impact performance but do not completely break the system. - Example: Improper training loop structure, incorrect data augmentation, or missing essential components in data processing. - Low: Errors in specific features that cause deviations from expected results but can be worked around with modifications. Any errors in the evaluation process belong to this category unless they impact the core concepts. These include minor issues like logging, error handling mechanisms, configuration settings, evaluation steps that do not alter the fundamental implementation and additional implementations not explicitly stated in the paper. - Example: Suboptimal hyperparameter initialization, incorrect learning rate schedule, inaccuracies in evaluation metrics, using a different random seed, variations in batch processing, different weight initialization, issues in result logging or reporting, variations in evaluation dataset splits, improper error handling in non-critical steps, mismatches in secondary evaluation criteria, or additional implementation details not specified in the paper that do not interfere with core results.
Prompt for model-based reference-free evaluation Example JSON format: “‘json { "critique_list": [ { "file_name": "dataset.py", "func_name": "train_preprocess", "severity_level": "medium", "critique": "A critique of the target repository’s file." }, { "file_name": "metrics.py", "func_name": "f1_at_k", "severity_level": "low", "critique": "A critique of the target repository’s file." } ], "score": 2 } “‘ Sample: Research Paper: {{The content of the paper}} Code Repository: {{The generated repository, which serves as the target repository for evaluation.}} Please provide a critique list for the code repository and a single numerical rating (1, 2, 3, 4, or 5) based on the quality of the sample, following the Example JSON format, without any additional commentary, formatting, or chattiness.
Figure 15: Prompts for model-based reference-free evaluation. {{}} indicate placeholders to be filled with the content described in the accompanying explanation.
Appendix F Examples output of the planning phase
Figure 16: Artifact from Step 1.1 (Overall Plan) in the planning stage of PaperCoder, generated during repository construction for the Transformer [42] (1).
Figure 17: Artifact from Step 1.1 (Overall Plan) in the planning stage of PaperCoder, generated during repository construction for the Transformer [42] (2).
Figure 18: Artifact from Step 1.2 (Architecture Design) in the planning stage of PaperCoder, generated during repository construction for the Transformer [42].
Figure 19: Class diagram visualization generated from Step 1.2 (Architecture Design) in the planning stage of PaperCoder, generated during repository construction for the Transformer [42].
Figure 20: Program call flow visualization generated from Step 1.2 (Architecture Design) in the planning stage of PaperCoder, generated during repository construction for the Transformer [42].
Figure 21: Artifact from Step 1.3 (Logic Design) in the planning stage of PaperCoder, generated during repository construction for the Transformer [42].
## config.yaml
training:
optimizer: "Adam"
beta1: 0.9
beta2: 0.98
epsilon: 1e-9
warmup_steps: 4000
total_steps: 100000 # Base model: 100K training steps (~12 hours on 8 P100 GPUs)
learning_rate_schedule: "d_model^-0.5 * min(step^-0.5, step * warmup_steps^-1.5)"
batch_tokens:
source: 25000
target: 25000
dropout: 0.1
label_smoothing: 0.1
model:
num_layers: 6
d_model: 512
d_ff: 2048
num_heads: 8
d_k: 64
d_v: 64
shared_embedding: true
data:
dataset: "WMT 2014 English-to-German"
encoding: "byte-pair encoding"
vocabulary_size: 37000
evaluation:
beam_size: 4
length_penalty: 0.6
max_output_length: "input_length + 50"
Figure 22: Artifact from Step 1.4 (Configuration File) in the planning stage of PaperCoder, generated during repository construction for the Transformer [42].
Table 11: List of ICLR 2024 papers used in our Paper2CodeBench benchmark. We evaluate each paper using the model-based, reference-free setting, with gpt-4o-2024-11-20 as the evaluation model.
Paper
Source Score
Generative Judge for Evaluating Alignment
Poster 4
Distributional Preference Learning: Understanding and Accounting for Hidden Context in RLHF
Poster 4
Inherently Interpretable Time Series Classification via Multiple Instance Learning
Oral 3.9
iTransformer: Inverted Transformers Are Effective for Time Series Forecasting
Oral 3.9
Tell Your Model Where to Attend: Post-hoc Attention Steering for LLMs
Poster 3.9
Knowledge Distillation Based on Transformed Teacher Matching
Poster 3.9
Meaning Representations from Trajectories in Autoregressive Models
Poster 3.8
A Simple Interpretable Transformer for Fine-Grained Image Classification and Analysis
Poster 3.8
VDC: Versatile Data Cleanser based on Visual-Linguistic Inconsistency by Multimodal Large Language Models
Poster 3.8
Vocos: Closing the gap between time-domain and Fourier-based neural vocoders for high-quality audio synthesis
Poster 3.8
SliceGPT: Compress Large Language Models by Deleting Rows and Columns
Poster 3.8
Beyond Accuracy: Evaluating Self-Consistency of Code Large Language Models with IdentityChain
Poster 3.8
Guiding Masked Representation Learning to Capture Spatio-Temporal Relationship of Electrocardiogram
Poster 3.8
Social Reward: Evaluating and Enhancing Generative AI through Million-User Feedback from an Online Creative Community
Oral 3.7
Language Model Detectors Are Easily Optimized Against
Poster 3.7
Improving protein optimization with smoothed fitness landscapes
Poster 3.7
SparseFormer: Sparse Visual Recognition via Limited Latent Tokens
Poster 3.7
AutoVP: An Automated Visual Prompting Framework and Benchmark
Poster 3.7
Hierarchical Context Merging: Better Long Context Understanding for Pre-trained LLMs
Poster 3.7
SEABO: A Simple Search-Based Method for Offline Imitation Learning
Poster 3.7
OpenChat: Advancing Open-source Language Models with Mixed-Quality Data
Poster 3.7
Rethinking The Uniformity Metric in Self-Supervised Learning
Poster 3.7
VONet: Unsupervised Video Object Learning With Parallel U-Net Attention and Object-wise Sequential VAE
Poster 3.6
Efficient Backpropagation with Variance-Controlled Adaptive Sampling
Poster 3.6
Structuring Representation Geometry with Rotationally Equivariant Contrastive Learning
Poster 3.6
ControlVideo: Training-free Controllable Text-to-Video Generation
Poster 3.6
Context-Aware Meta-Learning
Poster 3.6
RECOMBINER: Robust and Enhanced Compression with Bayesian Implicit Neural Representations
Poster 3.6
Peering Through Preferences: Unraveling Feedback Acquisition for Aligning Large Language Models
Poster 3.6
Modulate Your Spectrum in Self-Supervised Learning
Poster 3.6
Table 12: List of ICML 2024 papers used in our Paper2CodeBench benchmark. We evaluate each paper using the model-based, reference-free setting, with gpt-4o-2024-11-20 as the evaluation model.
Paper
Source Score
SAMformer: Unlocking the Potential of Transformers in Time Series Forecasting with Sharpness-Aware Minimization and Channel-Wise Attention
Oral 4
Autoformalizing Euclidean Geometry
Poster 4
Recurrent Distance Filtering for Graph Representation Learning
Poster 4
CosPGD: an efficient white-box adversarial attack for pixel-wise prediction tasks
Poster 3.9
Token-level Direct Preference Optimization
Poster 3.9
BayOTIDE: Bayesian Online Multivariate Time Series Imputation with Functional Decomposition
Oral 3.8
CurBench: Curriculum Learning Benchmark
Poster 3.8
Exploring the Low-Pass Filtering Behavior in Image Super-Resolution
Poster 3.8
Towards Efficient Exact Optimization of Language Model Alignment
Poster 3.7
On the Effectiveness of Supervision in Asymmetric Non-Contrastive Learning
Poster 3.7
Drug Discovery with Dynamic Goal-aware Fragments
Poster 3.7
Fool Your (Vision and) Language Model With Embarrassingly Simple Permutations
Poster 3.7
Image Restoration Through Generalized Ornstein-Uhlenbeck Bridge
Poster 3.7
Timer: Generative Pre-trained Transformers Are Large Time Series Models
Poster 3.7
Mitigating Oversmoothing Through Reverse Process of GNNs for Heterophilic Graphs
Poster 3.7
Scribble-Supervised Semantic Segmentation with Prototype-based Feature Augmentation
Poster 3.7
ConvNet vs Transformer, Supervised vs CLIP: Beyond ImageNet Accuracy
Poster 3.7
CLIF: Complementary Leaky Integrate-and-Fire Neuron for Spiking Neural Networks
Oral 3.6
FiT: Flexible Vision Transformer for Diffusion Model
Oral 3.6
Decomposing Uncertainty for Large Language Models through Input Clarification Ensembling
Oral 3.6
SparseTSF: Modeling Long-term Time Series Forecasting with *1k* Parameters
Oral 3.6
Sample-specific Masks for Visual Reprogramming-based Prompting
Oral 3.6
Boundary Exploration for Bayesian Optimization With Unknown Physical Constraints
Poster 3.6
Listwise Reward Estimation for Offline Preference-based Reinforcement Learning
Poster 3.6
Graph Distillation with Eigenbasis Matching
Poster 3.6
Temporal Spiking Neural Networks with Synaptic Delay for Graph Reasoning
Poster 3.6
Position: Quo Vadis, Unsupervised Time Series Anomaly Detection?
Poster 3.6
Neural SPH: Improved Neural Modeling of Lagrangian Fluid Dynamics
Poster 3.6
Self-Play Fine-Tuning Converts Weak Language Models to Strong Language Models
Poster 3.6
Unveiling and Harnessing Hidden Attention Sinks: Enhancing Large Language Models without Training through Attention Calibration
Poster 3.6
Table 13: List of NeurIPS 2024 papers used in our Paper2CodeBench benchmark. We evaluate each paper using the model-based, reference-free setting, with gpt-4o-2024-11-20 as the evaluation model.
Paper
Source Score
PACE: marrying generalization in PArameter-efficient fine-tuning with Consistency rEgularization
Oral 4
The Road Less Scheduled
Oral 4
G-Retriever: Retrieval-Augmented Generation for Textual Graph Understanding and Question Answering
Poster 4
Binarized Diffusion Model for Image Super-Resolution
Poster 4
Learning to Predict Structural Vibrations
Poster 4
Attack-Aware Noise Calibration for Differential Privacy
Poster 4
Make Your LLM Fully Utilize the Context
Poster 3.9
Smoothed Energy Guidance: Guiding Diffusion Models with Reduced Energy Curvature of Attention
Poster 3.9
Sm: enhanced localization in Multiple Instance Learning for medical imaging classification
Poster 3.9
AutoTimes: Autoregressive Time Series Forecasters via Large Language Models
Poster 3.9
End-to-End Ontology Learning with Large Language Models
Poster 3.8
Scaling transformer neural networks for skillful and reliable medium-range weather forecasting
Poster 3.8
Autoregressive Image Generation without Vector Quantization
Oral 3.7
Adaptive Randomized Smoothing: Certified Adversarial Robustness for Multi-Step Defences
Oral 3.7
Generalizable Person Re-identification via Balancing Alignment and Uniformity
Poster 3.7
Universal Neural Functionals
Poster 3.7
Are Self-Attentions Effective for Time Series Forecasting?
Poster 3.7
xMIL: Insightful Explanations for Multiple Instance Learning in Histopathology
Poster 3.7
Leveraging Environment Interaction for Automated PDDL Translation and Planning with Large Language Models
Poster 3.7
Task-Agnostic Machine Learning-Assisted Inference
Poster 3.7
Make Continual Learning Stronger via C-Flat
Poster 3.7
DARG: Dynamic Evaluation of Large Language Models via Adaptive Reasoning Graph
Poster 3.7
AsyncDiff: Parallelizing Diffusion Models by Asynchronous Denoising
Poster 3.7
You Only Look Around: Learning Illumination Invariant Feature for Low-light Object Detection
Poster 3.6
MutaPLM: Protein Language Modeling for Mutation Explanation and Engineering
Poster 3.6
Advancing Training Efficiency of Deep Spiking Neural Networks through Rate-based Backpropagation
Poster 3.6
Improved off-policy training of diffusion samplers
Poster 3.6
Navigating the Effect of Parametrization for Dimensionality Reduction
Poster 3.6
Long-Range Feedback Spiking Network Captures Dynamic and Static Representations of the Visual Cortex under Movie Stimuli
Poster 3.6
InfLLM: Training-Free Long-Context Extrapolation for LLMs with an Efficient Context Memory
Poster 3.6
Table 14: List of papers used in human evaluation. We evaluate the official repository of each paper, released by the authors, using the model-based reference-free setting with gpt-4o-2024-11-20 as the evaluation model.
RepoName
Paper
Score
VideoICL
VideoICL: Confidence-based Iterative In-context Learning for Out-of-Distribution Video Understanding
2.6
MuDI
Identity Decoupling for Multi-Subject Personalization of Text-to-Image Models
3.3
KALMV
Knowledge-Augmented Language Model Verification
3.3
sea-attention
SEA: Sparse Linear Attention with Estimated Attention Mask
2.7
HarmAug
HarmAug: Effective Data Augmentation for Knowledge Distillation of Safety Guard Models
3.0
GruM
Graph Generation with Diffusion Mixture
3.7
Adaptive-RAG
Adaptive-RAG: Learning to Adapt Retrieval-Augmented Large Language Models through Question Complexity
2.7
SoT
Sketch-of-Thought: Efficient LLM Reasoning with Adaptive Cognitive-Inspired Sketching
4.0
Mol-LLaMA
Mol-LLaMA: Towards General Understanding of Molecules in Large Molecular Language Model
3.5
judge_code_efficiency
Rethinking Code Refinement: Learning to Judge Code Efficiency
3.1
KARD
Knowledge-Augmented Reasoning Distillation for Small Language Models in Knowledge-Intensive Tasks
3.2
COINCIDE_code
Concept-skill Transferability-based Data Selection for Large Vision-Language Models
3.0
Janus
Aligning to thousands of preferences via system message generalization
3.5
N/A
Silent Branding Attack: Trigger-free Data Poisoning Attack on Text-to-Image Diffusion Models
N/A
VideoRAG
VideoRAG: Retrieval-Augmented Generation over Video Corpus
3.0
RADA
Retrieval-augmented data augmentation for low-resource domain tasks
3.0
STELLA_code
STELLA: Continual Audio-Video Pre-training with Spatio-Temporal Localized Alignment
3.3
prometheus-vision
Prometheus-vision: Vision-language model as a judge for fine-grained evaluation
3.1
CoLoR
Efficient Long Context Language Model Retrieval with Compression
3.0
Volcano
Volcano: Mitigating Multimodal Hallucination through Self-Feedback Guided Revision
3.2
N/A
T1: Tool-integrated Self-verification for Test-time Compute Scaling in Small Language Models
N/A
Table 15: List of papers used in executability analysis.
Repo Name
Paper
CoLoR
Efficient Long Context Language Model Retrieval with Compression
cognitive-behaviors
Cognitive Behaviors that Enable Self-Improving Reasoners, or, Four Habits of Highly Effective STaRs
RADA
Retrieval-Augmented Data Augmentation for Low-Resource Domain Tasks
Self-Instruct
Self-Instruct: Aligning Language Models with Self-Generated Instructions
G-EVAL
G-Eval: NLG Evaluation using GPT-4 with Better Human Alignment
Figure 23: Human Evaluation Format
Figure 24: Human Evaluation Format
Figure 25: Human Evaluation Format